### v1

# On the dynamics of leapfrogging vortex rings

Abstract.
The evolution of highly concentrated vorticity around rings in the three-dimensional axisymmetric Euler equations is studied in a regime for which the leapfrogging dynamics predicted by Helmholtz is expected to occur. For a general class of initial data corresponding to vortex rings of sufficiently small thickness $\varepsilon>0$, we prove weak and strong localization estimates for the vortex rings at positive times up to some time $T_{0}/|\ln\varepsilon|$ for fixed $T_{0}$, independent of $\varepsilon$. Moreover, we derive the motion law for the position of the rings on $[0,T_{0}/|\ln\varepsilon|]$.
To the best of our knowledge, this is the first result of a rigorous derivation of the leapfrogging motion up to times of order $|\ln\varepsilon|^{-1}$ for general initial data and the suitable scaling regime. The singular interaction of rings requires significant improvements of weak and strong localization estimates obtained in prior works. Our method is based on the combination of a new variational argument and a recently introduced double iterative procedure. Possibly due to natural filamentation of the vortex rings, which induces a lack of control of the support of the solution in the symmetry axis direction, our method does not allow to reach the time at which the rings have performed a full rotation around each other.
1. Introduction
As a model of the evolution of an incompressible and inviscid three-dimensional fluid, we consider the standard 3D Euler equations
(1.1) $\begin{cases}\partial_{t}\mathbf{U}+(\mathbf{U}\cdot\nabla)\mathbf{U}=-\nabla P,\\ \operatorname{div}\mathbf{U}=0,\end{cases}$
where for $(t,x)\in[0,T]\times{\mathbb{R}}^{3}$, the quantity $\mathbf{U}(t,x)$ denotes the velocity of the fluid at time $t$ at the position $x$. Letting $\mathbf{\Omega}=\nabla\wedge\mathbf{U}$ denote the vorticity of the fluid, (1.1) can be expressed as
(1.2) $\begin{cases}\partial_{t}\mathbf{\Omega}+(\mathbf{U}\cdot\nabla)\mathbf{\Omega}=(\mathbf{\Omega}\cdot\nabla)\mathbf{U}\vspace{2mm}\\ \displaystyle\mathbf{U}(t,x)=-\frac{1}{4\pi}\int_{{\mathbb{R}}^{3}}\frac{x-y}{|x-y|^{3}}\wedge\mathbf{\Omega}(t,y)\mathrm{d}y,\end{cases}$
where $\mathbf{U}$ is uniquely determined by $\mathbf{\Omega}$ for all $t\in[0,T]$ by means of the Biot-Savart law (the second equation in (1.2)). In the present paper, we restrict our attention to solutions that are axisymmetric vector fields without swirl, namely if considered in cylindrical coordinates $(e_{z},e_{r},e_{\theta})$, the solution $\mathbf{U}$ satisfies
$\mathbf{U}(z,r,\theta)=U_{z}(z,r)e_{z}+U_{r}(z,r)e_{r},$
i.e. $\mathbf{U}$ is independent of the toroidal variable $\theta$ (axisymmetric) and $\mathbf{U}\cdot e_{\theta}=0$ (without swirl). This implies that the vorticity $\mathbf{\Omega}=\nabla\wedge\mathbf{U}$ is a purely toroidal vector field:
$\mathbf{\Omega}=\Omega e_{\theta},\quad\text{with}\,\,\Omega(z,r)=\operatorname{curl}\begin{pmatrix}U_{z}\\ U_{r}\end{pmatrix}=(\partial_{z}U_{r}-\partial_{r}U_{z})(z,r).$
Problem (1.2) can hence be reduced to a two-dimensional problem for the potential scalar vorticity $\omega=r^{-1}\Omega$ posed on the half plane $(z,r)\in{\mathbb{R}}_{+}^{2}={\mathbb{R}}\times{\mathbb{R}}_{+}$, see Section 2 for details:
(1.3) $\begin{cases}\partial_{t}\,\omega+u\cdot\nabla\omega=0&\text{ in }{\mathbb{R}}_{+}^{*}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{div}(ru)=0&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{curl}(u)=r\omega&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \omega(0,\cdot)=\omega_{0}&\text{ in }{\mathbb{R}}^{2}_{+}.\end{cases}$
The scope of this paper is to investigate the dynamics when the vorticity is sharply concentrated around rings. Specifically, we consider initially concentrated vortex rings of thickness $\varepsilon$, for sufficiently small $\varepsilon$ and we rigorously identify the motion of rings at leading order. The study of such vortex rings goes back to Helmholtz [29] and J. J. Thomson [48]. This is the subject of many studies ranging from mathematics to engineering. They are observed in 3D flows with high Reynolds number as long-lived vortex structures that translate along their symmetry axis. It was computed by Kelvin [36] that a vortex ring of major radius $r$, minor radius $\varepsilon$, and circulation $\gamma$ has a self-induced velocity
(1.4) $v=\frac{\gamma}{4\pi r}|\ln\varepsilon|e_{z},$
at the first order in $\varepsilon$, see also [25, 5] for rigorous derivation. The respective motion law for a single vortex ring was rigorously established for weakly viscous fluids in [7, 9, 27]. More generally, assuming that the vorticity is concentrated around a curve, the motion is expected to be asymptotically governed by the binormal curvature flow equation, as was formally computed by Da Rios [47]. The binormal curvature flow exhibits a rich panel of dynamics that have been intensively studied. In particular a serie of papers by Banica and Vega is devoted to self-similar solutions or infinite polygonal lines solutions (we refer to [4] for a review of the results). Finally, Jerrard and Smets [34] introduced a new formulation of the binormal curvature flow equation allowing for weaker solutions.
The issue of the persistence in time of such vortex filaments by the Euler flow is the so-called vortex filament conjecture, see [33, 24] for partial results in this direction. While the conjecture remains open in its generality, results are available for specific geometric and dimension reduced settings. The rigorous asymptotic dynamics of 2D point vortices (the motion of Dirac masses of vorticity in the 2D Euler equations) was obtained in [40]. In the aforementioned case of annular vortex rings evolving according to (1.3), positive answers have been given with various assumptions. We refer to the work by Buttà, Cavallaro and Marchioro [8, 11] and references therein. This is in particular due to the fact that the problem allows for a dimension reduction leading to the 2D problem (1.3).
We also mention that the desingularization problem for the vortex rings has been addressed recently by Cao, Wan and Zhan [13] and by Davila, Del Pino, Musso and Wei [18].
We emphasize that the system (1.3) notably differs from the 2D Euler equations. Indeed, the dynamics of 2D Euler point-vortices do not include any self-induced motion of vortices. Even though the problem (1.3) is reduced to two dimensions, as in the study of helical filaments in [20] or point-vortices for the lake equations [31], the Biot-Savart law (the relation giving $u$ in terms of $\omega$) is much more complex compared to the planar case. Additional techniques are required and the results developed for the 2D Euler equations (see for instance [40]) cannot be applied. Roughly speaking, the motion of vortex rings is driven by the self-induced motion (1.4) and by their mutual interactions. Depending on the chosen asymptotic regime, the former maybe dominant, leading to a translation of the vortex rings along the symmetry axis [12, 8]. In case the latter dominates, the dynamics is reminiscent to the 2D dynamics of point vortices [10]. In particular, it was shown in [8] that the interactions of vortex rings of different radii (independent of $\varepsilon$) are bounded independently of $\varepsilon$ and can be neglected compared to their self-interaction (inducing the velocity given by (1.4)) which is of order $|\ln\varepsilon|$.
However, it is well-known since the work of Helmholtz [29], see also [21, 22, 30, 37, 6], and it has been confirmed by numerical [3, 38, 45] and experimental evidence [51] that in particular conditions, two coaxial vortex rings can successively pass through each other in a periodic motion known as the leapfrogging phenomenon. For such phenomena to occur provided, the vortex rings need to be located at infinitesimal mutual distance leading to mutual interactions inducing a change of their radii that is significant enough to cause a variation of their self-induced velocity (1.4). We say that leapfrogging dynamics occurs when the variation of vertical velocity in the self-interaction is of the same order as the mutual interactions of ring. It turns out that this happens when the vortex rings are placed a distance of order $1/\sqrt{|\ln\varepsilon|}$ to each other, and a detailed explanation of the suitable scaling regime is provided in Appendix B, see also [18, 35]. In the recent paper [18], the authors construct a smooth solution to (1.3) exhibiting the leapfrogging motion. The 3D Gross-Pitaevskii equation also exhibits leapfrogging dynamics as proven by Jerrard and Smets [35].
In the present paper, we rigorously derive the leapfrogging dynamics for (1.3) in the aforementioned scaling regime and a general class of initial data that remained open despite being expected since [29], see Theorem 1.1 below. We point out that the present scaling regime differs from the one in [11] in which the vortex positions are located at a mutual distance of order $1/|\ln\varepsilon|$. While the mutual interactions are more singular compared to our setting, the estimates obtained there hold on a significantly shorter time interval. The differences between the regimes are discussed in detail in Appendix B.
In the present setting, all vortex rings move at leading order at the same vertical velocity $v\propto|\ln\varepsilon|$ (relation (1.4)), at a relative distance of order $1/\sqrt{|\ln\varepsilon|}$ and in a relative motion of velocity $v^{\prime}\propto\sqrt{|\ln\varepsilon|}$. Therefore, it is very natural to rescale the time variable by a factor $|\ln\varepsilon|$, so that both motions happen on a time of order 1. More precisely, we introduce the system
(1.5) $\begin{cases}\displaystyle\partial_{t}\,\omega^{\varepsilon}+\frac{1}{|\ln\varepsilon|}u^{\varepsilon}\cdot\nabla\omega^{\varepsilon}=0&\text{ in }{\mathbb{R}}_{+}^{*}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{div}u^{\varepsilon}=0&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{curl}(u^{\varepsilon})=r\omega^{\varepsilon}&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \omega^{\varepsilon}(0,\cdot)=\omega_{0}^{\varepsilon}&\text{ in }{\mathbb{R}}^{2}_{+}.\end{cases}$
for which we expect two vortex rings at a distance of order $1/\sqrt{|\ln\varepsilon|}$ would perform a complete rotation around each other in a time of order 1, while traveling altogether a vertical distance also of order 1.
We can now formulate our problem in a precise mathematical framework and state our result. For $x\in{\mathbb{R}}^{2}_{+}$, we further denote by $x=(x_{z},x_{r})$ with $x_{z}\in{\mathbb{R}}$ and $x_{r}\in{\mathbb{R}}_{+}$, and by $(e_{z},e_{r})$ the canonical basis on ${\mathbb{R}}^{2}$. We consider a fixed reference point:
$X^{*}=(z^{\ast},r^{\ast}).$
For $\varepsilon>0$, the initial data $\omega_{0}^{\varepsilon}$ will be a superposition of $N$ vorticities sharply concentrated around points located at distance of order $1/\sqrt{|\ln\varepsilon|}$ from the reference point:
(1.6) $X^{\varepsilon}_{i,0}:=X^{*}+\frac{Y_{i,0}}{\sqrt{|\ln\varepsilon|}},$
where $Y_{i,0}\in{\mathbb{R}}^{2}$ is fixed independently of $\varepsilon$. We next assume that the vortex filaments all have the same density $\gamma$. Namely, let $\gamma\neq 0$, we assume that $\omega_{0}^{\varepsilon}\in L^{\infty}({\mathbb{R}}^{2}_{+})$ such that
(1.7) $&\omega_{0}^{\varepsilon}=\sum_{i=1}^{N}\omega_{i,0}^{\varepsilon},\\ &\operatorname{supp}\omega_{i,0}^{\varepsilon}\subset B\big{(}X_{i,0}^{\varepsilon},\varepsilon\big{)},\\ &\omega_{i,0}^{\varepsilon}\text{ has a definite sign and }\int_{{\mathbb{R}}^{2}_{+}}x_{r}\omega_{i,0}^{\varepsilon}(x)\mathrm{d}x=\gamma,\\ &|\omega_{0}^{\varepsilon}|\leqslant\frac{M_{0}}{\varepsilon^{2}},\end{aligned}\right.$
for some $M_{0}>0$ independent of $\varepsilon$. The goal of this paper is to prove that the sequence of vorticities $\omega^{\varepsilon}(t,\cdot)$, solution of (1.5) for such intial data, remains asymptotically concentrated around $N$ points $X_{i}^{\varepsilon}(t)$ and that the trajectories of these points is governed by the ODE:
(1.8) $\begin{cases}\displaystyle\frac{\mathrm{d}}{\mathrm{d}t}X_{i}^{\varepsilon}(t)=\frac{\gamma}{4\pi X_{i,r}^{\varepsilon}(t)}e_{z}+\frac{\gamma}{2\pi|\ln\varepsilon|}\sum_{j\neq i}\frac{\big{(}X_{i}^{\varepsilon}(t)-X_{j}^{\varepsilon}(t)\big{)}^{\perp}}{\big{|}X_{i}^{\varepsilon}(t)-X_{j}^{\varepsilon}(t)\big{|}^{2}},\\ X_{i}^{\varepsilon}(0)=X_{i,0}^{\varepsilon}.\end{cases}$
Before stating our result rigorously, we briefly comment on the ODE (1.8). We recovered, in the first term, the vertical self-induced velocity (1.4), and in the second term, the well-known mutual interaction in the 2D point-vortex dynamics, both up to a factor $1/|\ln\varepsilon|$ coming from the time rescaling. The system (1.8) is clearly Hamiltonian. It is not known whether the solution is global, should several vortex rings collapse together at $r=0$. In Section 5.2, we prove the existence of a time $T_{X}$ depending on $(Y_{i,0})_{1\leqslant i\leqslant N}$ but not on $\varepsilon$ such that the system (1.8) has a solution on the time interval $[0,T_{X}]$. To this end, we observe that $t\mapsto Y_{i}^{\varepsilon}(t)$ defined by the relation
(1.9) $X_{i}^{\varepsilon}(t)=X^{*}+\frac{\gamma}{4\pi r^{*}}te_{z}+\frac{1}{\sqrt{|\ln\varepsilon|}}Y_{i}^{\varepsilon}(t),$
converge as $\varepsilon\to 0$ to a limiting trajectory $t\mapsto\widetilde{Y_{i}}(t)$. This limiting trajectory satisfies
[…]
and in works such as [35, 18, 42], it is this dynamics that is considered as the motion of the rings in the right scaling. For (1.10), no collisions occur in finite time, see Appendix C. In this paper we prefer to work with the trajectories $(X_{i})_{1\leqslant i\leqslant N}$ and $(Y_{i})_{1\leqslant i\leqslant N}$ as it is much more natural in the computations, then prove in Section 5 that in all the following, one could replace the trajectories $t\mapsto X_{i}^{\varepsilon}(t)$ by
(1.11) $\widetilde{X}_{i}^{\varepsilon}(t):=X^{*}+\frac{\gamma}{4\pi r^{*}}te_{z}+\frac{1}{\sqrt{|\ln\varepsilon|}}\widetilde{Y}_{i}(t)$
to match the notations of [35, 18, 42], and all the results, in particular the following Theorem 1.1, would still hold.
We are now in position to give precisely our main theorem.
Theorem 1.1.
Let $X^{*}=(z^{*},r^{*})\in{\mathbb{R}}^{2}_{+}$, $(Y_{i,0})_{1\leqslant i\leqslant N}$ be $N$ points in ${\mathbb{R}}^{2}$ such that $Y_{i,0,r}\neq Y_{j,0,r}$ for all $i\neq j$. Let $\gamma\neq 0$, $M_{0}>0$, and $\omega_{0}^{\varepsilon}\in L^{\infty}({\mathbb{R}}^{2}_{+})$ satisfying (1.7) for every $\varepsilon>0$.
There exists $T_{0}>0$, $\varepsilon_{0}>0$ and $C_{0}>0$, depending only on $X^{*}$, $(Y_{i,0})$ and $M_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, there exists a unique solution $(t\mapsto X_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$ of (1.8) on $[0,T_{0}]$, and the unique weak solution $(u^{\varepsilon},\omega^{\varepsilon})$ of (1.5) on $[0,T_{0}]$ (in the sense of Proposition  2.1 below) can be decomposed as
$\omega^{\varepsilon}=\sum_{i=1}^{N}\omega_{i}^{\varepsilon},\quad\omega_{i}^{\varepsilon}\in L^{\infty}({\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+})$
with the following properties:
  1. (i)
    A weak localization property:
    $\sup_{t\in[0,T_{0}]}\left|\gamma-\int_{B\left(X_{i}^{\varepsilon}(t),C_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x\right|\leqslant\gamma\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
  2. (ii)
    A strong localization property:
    $\operatorname{supp}\omega_{i}^{\varepsilon}(t,\cdot)\subset\left\{(z,r)\in{\mathbb{R}}^{2}_{+}\,,\,\big{|}r-X_{i,0,r}^{\varepsilon}\big{|}\leqslant\frac{1}{4}\frac{\min_{j\neq k}|Y_{j,0,r}-Y_{k,0,r}|}{\sqrt{|\ln\varepsilon|}}\,,\,\big{|}z-z^{*}\big{|}\leqslant 1\right\},\ \text{for all}\ t\in[0,T_{0}].$
Let us now formulate several important remarks regarding Theorem 1.1. Concerning previous results and to the best of our knowledge, the only rigorous results concerning the leapfrogging dynamics for the 3D Euler equations are [18, 11] and [35] for the 3D Gross-Pitaevskii equation. While [18, 35] consider the same scaling regime, the regime studied in [11] differs from the present one, as previously mentioned. The results from [11] cannot be adapted to our situation, as also pointed out by the authors in [11, p.791]. In [18], the authors construct initial data $\omega_{0}^{\varepsilon}$ of a very specific class to obtain a smooth solution exhibiting leapfrogging vortex rings. We wish to emphasize the fact that the conclusions of Theorem 1.1 hold for any initial data $\omega_{0}^{\varepsilon}\in L^{\infty}({\mathbb{R}}^{2}_{+})$ satisfying relations (1.7), which ranges from smooth compactly supported blobs to vortex patches, even possibly multiply connected. Opposite to [18], our proof does not rely on sophisticated elliptic desingularization techniques. The method in [35] is based on Jacobian estimates not available for (1.1).
The weak localization property, namely item $(i)$ in Theorem 1.1, is obtained by a new and original argument. Assuming weak and strong localization, we obtain a precise expansion of the vortex rings’ energy from which we deduce a concentration property, inspired by the work of Turkington in [49]. By doing so, we largely improve the previously known weak localization estimates [8]. Proposition 4.1 below provides even more precise estimates than what is stated in Theorem 1.1. In particular, for any $\eta>0$ it gives the existence of a domain of size $\varepsilon^{1-\eta}$ containing most of the potential vorticity. However, interestingly enough, our method fails to provide a precise location for this domain, see Remark 4.13 for a detailed discussion. Finally, a bootstrap argument yields that the weak concentration holds as long as the strong localization does.
The strong localization property $(ii)$ in Theorem 1.1 is pivotal for deriving the asymptotic motion and constitutes a substantial improvement compared to previously known results in [8] and similarly in [31, 20] in different geometric settings. These previous results state that the potential vorticity remains supported in a strip of radial size $1/|\ln\varepsilon|^{\kappa}$, for $\kappa<1/4$. Here, we prove a localization of radial size $1/\sqrt{|\ln\varepsilon|}$. This constitutes the critical size in order to observe a motion on spatial scales of order $1/\sqrt{|\ln\varepsilon|}$ on which leapfrogging occurs. This is achieved in two independent steps. We first prove the confinement in the radial direction by adapting the double iterative procedure recently introduced in [11] to the present regime. Second, we prove the localization in the vertical direction in a strip of size of order $1$ being less precise than the radial direction and therefore easier to obtain. On a related note, in the case of a single vortex ring, the result of Theorem 1.1 also improves those of [8].
In the following Remark 1.2, the localization property and its link to the small-time assumption for our main result are discussed.
Remark 1.2.
A natural and relevant question regarding our main result Theorem 1.1 is whether the small time assumption can be removed as it prevents to describe a full rotation of the vortex rings around each other. In [18], a particular smooth solution to (1.3) exhibiting the leapfrogging dynamics for an arbitrarily large number of rotations is constructed. In [11], an arbitrarily large number of rotations of two vortex rings around each other is observed provided that the vortex rings radii’ are large enough. As detailed in Appendix B, translating the result of [11] to the present setting by means of the scaling invariance of (1.3) yields a time scale of order $|\ln\varepsilon|^{-2}$ being much shorter than the present one.
Concerning possible generalization of our result, we expect that our small time assumption is not sharp and can be optimized. However, it becomes clear from our proof that our method is not expected to reach a full rotation. We only obtain a $\mathcal{O}(1)$ localization in the $z$ direction, and the strips in which each ring is supported are not allowed to overlap. Moreover, additional assumptions on the general class of initial data (1.7) considered here might be necessary. Indeed, it is conceivable that the possible occurrence of filamentation of the vortex ring in the vertical direction prevents an improvement of the $\mathcal{O}(1)$ localization in the $z$ direction for general initial data. Namely, a tail in the wake of the ring may appear, a phenomenon that was described for instance in [43, 44] and recently proved in [16] for Hill’s vortex. This fact may explain similar restrictions in [8, 12, 20, 31]. In [18], this filamentation is not present since the authors work with specific initial data and a stable vorticity profile is constructed. In [11], this filamentation does also not occur suggesting that the tail does appear on the short time scale considered.
Nevertheless, it might be possible to remove the small time assumption at least for the weak localization: this would require the weak localization to be obtained without assuming a priori strong localization, which appears not to be withing reach of our current method. For these reasons, the question of the stability of a concentrated vortex when colliding with the tail of another vortex seems to be an important and challenging direction for future research, which will require new tools.
An interesting quantity to estimate is the center of potential vorticity of each vortex blob in the 2D reduction, namely the expression
(1.12) $b^{\varepsilon}_{i}(t)=\int_{{\mathbb{R}}^{2}_{+}}x\frac{x_{r}\omega^{\varepsilon}_{i}(t,x)}{\gamma}\mathrm{d}x.$
We obtain in Proposition 5.4 as a consequence of the localization properties that
$\sup_{t\in[0,T_{0}]}\left|b_{i}^{\varepsilon}(t)-X_{i}^{\varepsilon}(t)\right|\leqslant C_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
Let us also mention that the strong localization $(ii)$ is obtained in a strip centered in $X_{i,0,r}$ for the sake of simplicity, one can easily adapt the proof to center the strip on $X_{i,r}(t)$ instead by actually centering the study on $b_{i,r}^{\varepsilon}(t)$.
The plan of the paper is the following. In Section 2, we recall well-posedness results for (1.3) and provide a suitable expansion for the Biot-Savart law. In Section 3, we introduce useful notations and gather properties on the interactions of vortex rings. In Section 4, we obtain the weak localization property by new bootstrap argument and assuming the strong localization, in an unspecified region of space. In Section 5, we obtain information on the motion of the center of mass and obtain the weak localization Theorem 1.1-$(i)$. In Section 6, we conclude the proof of Theorem 1.1 by proving the strong localization, successively in the $r$ and $z$ direction.
2. Well-posedness and Biot-Savart law
This section recalls and establishes well-posedness results for the axisymmetric 3D Euler equations without swirl.
2.1. Reduction
We denote by $(z,r,\theta)$ the usual cylindrical coordinates in ${\mathbb{R}}^{3}$ with basis vectors $(e_{z},e_{r},e_{\theta})$. A vector field $U:{\mathbb{R}}^{3}\rightarrow{\mathbb{R}}^{3}$ is called axisymmetric without swirl if of the form
$U(z,r,\theta)=U_{z}(z,r)e_{z}+U_{r}(z,r)e_{r},$
i.e. $U$ is independent of the toroidal variable $\theta$ (axisymmetric) and $U\cdot e_{\theta}=0$ (without swirl). This implies that the vorticity $\mathbf{\Omega}=\nabla\wedge U$ is a purely toroidal vector field:
$\mathbf{\Omega}=\Omega e_{\theta},\quad\text{with}\,\,\Omega=\partial_{z}U_{r}-\partial_{r}U_{z}.$
Here and below we adopt the notation $x\in{\mathbb{R}}_{+}^{2}={\mathbb{R}}\times(0,\infty)$ with $x=(x_{z},x_{r})$. Note the change of convention w.r.t. the usual cylindrical coordinates. This allows one to reduce (1.1) to a 2D problem by defining the scalar $\omega:{\mathbb{R}}_{+}^{2}\rightarrow{\mathbb{R}}$ and vector field $u:{\mathbb{R}}_{+}^{2}\rightarrow{\mathbb{R}}^{2}$ respectively as
$\omega=\frac{\Omega}{x_{r}},\qquad u=(U_{z},U_{r}).$
Note that the div-curl problem $\operatorname{div}U=0$ and $\nabla\wedge U=\mathbf{\Omega}$ reduces to
$\partial_{r}u_{r}+\frac{1}{x_{r}}u_{r}+\partial_{z}u_{z}=0,\quad x_{r}\omega=\partial_{z}u_{r}-\partial_{r}u_{z}.$
In particular, the elliptic problem posed on ${\mathbb{R}}_{+}^{2}={\mathbb{R}}\times(0,\infty)$ also writes
(2.1) $\operatorname{div}_{(z,r)}(x_{r}u)=0,\quad\operatorname{curl}_{(z,r)}(u)=x_{r}\omega.$
Under suitable boundary and regularity conditions, the velocity field $u$ is hence uniquely determined by $\omega$, see Section 2.3 for details. On the half plane ${\mathbb{R}}_{+}^{2}$, the potential vorticity $\omega$ satisfies the transport type equation
(2.2) $\partial_{t}\omega+u\cdot\nabla\omega=0,$
that also reads as continuity equation
(2.3) $\partial_{t}(x_{r}\omega)+\operatorname{div}(x_{r}u\omega)=0,$
given that $\operatorname{div}(x_{r}u)=0$ from (2.1).
2.2. Well-posedness results
Global existence and uniqueness of weak axisymmetric without swirl solutions of (1.1) has been proved in [50] and [46, Th 3.3], [1, 17] as well as references therein. In [14, Th. 5.4], global existence and uniqueness of the weak vorticity solution of (2.2) is proved when the initial velocity is axisymmetric without swirl, the initial vorticity satisfies $\Omega_{0}\in L^{1}\cup L^{\infty}({\mathbb{R}}_{+}^{2})$ and its support does not overlap the symmetry axis. More precisely, this solution satisfies
$\frac{\Omega(t,X(t,0,x))}{X_{r}(t,x)}=\frac{\Omega_{0}(x)}{x_{r}}$
where the flow map $X:[0,\infty)\times{\mathbb{R}}^{3}\to{\mathbb{R}}^{3}$ is defined by
$\frac{\mathrm{d}X(t,t_{0},x)}{\mathrm{d}t}=U\left(t,X(t,t_{0},x)\right),\quad X(t_{0},t_{0},x)=x.$
From now on we shall write $X(t,x)=X(t,0,x)$ for simplicity.
Since the velocity is axisymmetric without swirl, the flow map $X$ may be seen as a map on +×${\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+}$ to +${\mathbb{R}}^{2}_{+}$ namely $X=X_{z}(t,x_{z},x_{r})e_{z}+X_{r}(t,x_{z},x_{r})e_{z}$, where
(2.4) $\begin{split}\frac{\mathrm{d}X_{z}(t,x_{z},x_{r})}{\mathrm{d}t}&=u_{z}\left(t,X(t,x_{z},x_{r})\right),\quad X_{z}(0,x_{z},x_{r})=x_{z},\\ \frac{\mathrm{d}X_{r}(t,x_{z},x_{r})}{\mathrm{d}t}&=u_{r}\left(t,X(t,x_{z},x_{r}))\right),\quad X_{r}(0,x_{z},x_{r})=x_{r}.\end{split}$
In the solution constructed by [14, Th. 5.4], the velocity is reconstructed from the vorticity with the Biot-Savart Kernel: $U(t,x)=\int H(x-y)\wedge\mathbf{\Omega}(t,y)\,dy$ with $H(x)=-x/(4\pi|x|^{3})$ and with $\mathbf{\Omega}$ depending only on $(z,r)$ and purely toroidal. As used in [8, 12], the Biot-Savart law in this case admits an explicit representation:
[…]
Moreover, it is classical that $u$ is bounded (see (2.6) below) and log-lipschitz (by adapting, for instance, the planar case done in [41, App. 2.3]). Finally, it is established in [14, Corollary 5.6] that the solution is a weak solution of (2.2) in the sense of distributions (see (5.3) in [14]). We summarize these results on the half plane ${\mathbb{R}}_{+}^{2}$ in the next proposition.
Proposition 2.1 ([14, 27]).
Let $\Omega_{0}\in L^{1}\cap L^{\infty}({\mathbb{R}}_{+}^{2})$ and assume that $\inf_{x\in\operatorname{supp}(\Omega_{0})}x_{r}=\delta_{0}>0$. Let $\omega_{0}=\Omega_{0}/x_{r}$ so that $\omega_{0}\in L^{1}\cap L^{\infty}({\mathbb{R}}_{+}^{2})$. For all $T>0$ there exists a unique pair $(\omega,u)$ with $x_{r}\omega\in L^{\infty}([0,T];L^{1}\cap L^{\infty}({\mathbb{R}}_{+}^{2}))$, $x_{r}^{2}\omega\in L^{\infty}([0,T];L^{1}({\mathbb{R}}_{+}^{2}))$, $\omega\in L^{\infty}([0,T];L^{\infty}({\mathbb{R}}_{+}^{2}))$, and $u\in L^{\infty}([0,T]\times{\mathbb{R}}_{+}^{2})$, solution to (1.3) in the following sense
  1. (i)
    $u=(u_{z},u_{r})$ is given in terms of $\omega$ by the axisymmetric Biot-Savart law (2.5).
  2. (ii)
    The vorticity equation is satisfied in the sense of distribution, i.e. for all $\Phi\in C^{\infty}([0,T]\times{\mathbb{R}}_{+}^{2})$ and for any $t\in[0,T]$
    $\int_{{\mathbb{R}}_{+}^{2}}\Phi(t,x)x_{r}\omega(t,x)\mathrm{d}x-\int_{{\mathbb{R}}_{+}^{2}}\Phi(0,x)x_{r}\omega_{0}(x)\mathrm{d}x=\int_{0}^{t}\int_{{\mathbb{R}}_{+}^{2}}x_{r}\omega(s,x)\left(\partial_{t}\Phi+u\cdot\nabla\Phi\right)(s,x)\mathrm{d}x\mathrm{d}s.$
  3. (iii)
    We have $\omega(t,X(t,x))=\omega_{0}(x)$, where $X$ is the flow of $u$ defined in (2.4) [^1] Such a flow exists in the classical sense by log-Lipschitz regularity of $u$..
Moreover, we have the estimates
(2.6) $\|u\|_{L^{\infty}({\mathbb{R}}_{+}^{2})}\leqslant C\|x_{r}^{2}\omega\|_{L^{1}({\mathbb{R}}_{+}^{2})}^{1/2}\|\omega\|_{L^{\infty}({\mathbb{R}}_{+}^{2})}^{1/2}$
and
$\left|u(t,x)-u(t,y)\right|\leqslant C(\|x_{r}\omega\|_{L^{1}\cap L^{\infty}([0,T]\times{\mathbb{R}}_{+}^{2})})|x-y|\left(1+\left|\ln|x-y|\right|\right).$
A direct consequence of the log-Lipschitz estimate of the velocity field is the following:
Corollary 2.2.
Let $\omega_{0}\in L^{1}\cap L^{\infty}({\mathbb{R}}_{+}^{2})$ such that $x_{r}\omega_{0}\in L^{1}\cap L^{\infty}({\mathbb{R}}_{+}^{2})$ be as in Proposition 2.1 above. Let $T>0$ and $(\omega,u)$ be the unique weak solution of (1.3) on $[0,T]$ provided by Proposition 2.1. Then the following holds true.
  1. (i)
    There exists $\delta_{T}>0$ depending only on $\|\omega_{0}\|_{L^{\infty}}$, $\delta_{0}$ and $T$ such that
    $\operatorname{supp}\omega(t,\cdot)\subset\{x\,|\,x_{r}\geqslant\delta_{T}\},\quad\forall t\in[0,T].$
  2. (ii)
    If $\omega_{0}$ has compact support in ${\mathbb{R}}_{+}^{2}$, denoting by $R_{0}=\sup\{|x|\,|\,x\in\operatorname{supp}(\omega_{0})\}$, there exists a compact set $K_{T}\subset{\mathbb{R}}_{+}^{2}$ depending only on $\|x_{r}\omega_{0}\|_{L^{\infty}}$, $\delta_{0},R_{0}$ and $T$ such that
    $\operatorname{supp}\omega(t,\cdot)\subset K_{T},\quad\forall t\in[0,T].$
  3. (iii)
    For all $t\in[0,T]$ and $p\in[1,\infty)$ it holds $\|x_{r}^{\frac{1}{p}}\omega(t,\cdot)\|_{L^{p}}=\|x_{r}^{\frac{1}{p}}\omega_{0}\|_{L^{p}}$.
Proof.
We provide a quick proof for sake of completeness although this is probably a standard fact. Since $\omega$ is constant along trajectories it suffices to provide estimates on the Lagrangian flow. Let $x\in\operatorname{supp}(\omega_{0})$. Writing $X(t)=X(t,x)$ and observing that $u_{r}(t,x_{z},0)=0$ in view of (2.5) we have
$\left|\frac{\mathrm{d}X_{r}(t)}{\mathrm{d}t}\right|$ $=\left|u_{r}\left(t,X(t)\right)\right|=\left|u_{r}\left(t,X_{z}(t),X_{r}(t)\right)-u_{r}\left(t,X_{z}(t),0\right)\right|$
$\leqslant C|X_{r}(t)||\ln|X_{r}(t)||$
where we have used the log-Lipschitz property for $u$. Hence this proves $(i)$ by means of a Gronwall-type argument (see e.g. the arguments to obtain (6.26) from (6.23) in [41]). Property $(ii)$ is a straightforward consequence of the $L^{\infty}$ bound on the velocity. Finally, $(iii)$ follows from the continuity equation (2.3). ∎
Remark 2.3.
For initial vorticity as in (1.7) consisting of $N$ disjoint blobs, the existence of a unique Lagrangian flow allows one to decompose the unique weak solution as
(2.7) $\omega^{\varepsilon}(t,x)=\sum_{i=1}^{N}\omega_{i}^{\varepsilon}(t,x),\qquad\omega_{i}^{\varepsilon}(t,X(t,\cdot))=\omega_{i,0}^{\varepsilon}.$
Moreover, Part (i) of Corollary 2.2 states that there exists a positive time such that a vorticity that is initially localized away from the symmetry axis does not reach the axis and remains localized. Concerning the localization in the radial direction, a quantitative estimate of this type is stated in [15, Section 2.3], it is shown that the radius of the support
$R_{\varepsilon}(t)=\sup\{x_{r}:x\in\operatorname{supp}(\omega^{\varepsilon}(t,\cdot))\}$
grows at most as
$R_{\varepsilon}(t)\leqslant C_{\varepsilon}(1+t)^{2},$
where $C_{\varepsilon}=C(\|\omega_{0}^{\varepsilon}\|_{L^{1}({\mathbb{R}}^{3})},\|\omega_{0}^{\varepsilon}\|_{L^{\infty}({\mathbb{R}}^{3})})>0$. Unfortunately, these estimates of the radius and of the distance to the axis depend on $\|\omega_{0}^{\varepsilon}\|_{L^{\infty}({\mathbb{R}}^{3})}$ which blows up as $\varepsilon^{-2}$ when $\varepsilon\to 0$. Hence, we will not be able to use this kind of estimate and we will later define a time $T_{\varepsilon}$ such that the support of $\omega^{\varepsilon}$ is included in a strip ×${\mathbb{R}}\times[r^{*}/2,2r^{*}]$. One of the main objectives is to get a strong localization property to prove that $T_{\varepsilon}\geqslant T_{0}$ for all $\varepsilon\in(0,\varepsilon_{0}]$, provided that $T_{0}$ and $\varepsilon_{0}$ are chosen small enough.
2.3. Expansion of the Biot-Savart law
In the 3D axisymmetric framework, the Biot-Savart law admits an explicit representation formula (2.5). A suitable expansion of the Biot-Savart kernel used in sequel is introduced. We notice that the 2D div-curl problem (2.1) including the weighted incompressibility condition yields that there exists a stream function $\Psi:{\mathbb{R}}_{+}^{2}\rightarrow{\mathbb{R}}$ such that
(2.8) $u=\begin{pmatrix}u_{z}\\ u_{r}\end{pmatrix}=\frac{1}{x_{r}}\nabla^{\perp}\Psi=\frac{1}{x_{r}}\begin{pmatrix}-\partial_{r}\Psi\\ \partial_{z}\Psi\end{pmatrix}.$
Following [23, 26] the boundary conditions $\Psi(x_{z},0)=0$ are imposed. We are led to solve
(2.9) $\begin{cases}\displaystyle\operatorname{div}\left(\frac{1}{x_{r}}\nabla\Psi(x)\right)=x_{r}\omega(x)\\ \Psi(x_{z},0)=0.\end{cases}$
In relation with (2.5), there is an explicit formula of the Green kernel, namely it is given in [26, Section 2] that
(2.10) $\Psi(x)=\int_{{\mathbb{R}}_{+}^{2}}\mathcal{G}(x,y)y_{r}\omega(y)\mathrm{d}y,$
where
(2.11) $\mathcal{G}(x,y)=-\frac{1}{2\pi}\sqrt{x_{r}y_{r}}F\left(\frac{|x-y|^{2}}{x_{r}y_{r}}\right)\quad\text{with}\quad F(s)=\int_{0}^{\frac{\pi}{2}}\frac{\cos(2\theta)}{\sqrt{\sin^{2}(\theta)+\frac{s}{4}}}\mathrm{d}\theta,\quad s>0.$
This function $F$ was intensively studied in the literature. We bring together in the next lemma some estimates we need, and we refer to [26, Lemma 2.1] and [23, Lemma 2.7] for more details.
Lemma 2.4.
For any $R_{0}>0$ given, there exists $C_{0}>0$ such that the following estimates holds for any $s\in(0,R_{0}]$:
$\max\Bigg{(}\Big{|}F(s)+\frac{1}{2}\ln(s)\Big{|},\Big{|}F^{\prime}(s)+\frac{1}{2s}\Big{|},|s^{2}F^{\prime\prime}(s)|\Bigg{)}\leqslant C_{0}.$
The explicit formula of the leading term of $F$ allows for a decomposition of the Green function $\mathcal{G}$.
Corollary 2.5.
We have the following decomposition
$\mathcal{G}(x,y)=\frac{1}{2\pi}\sqrt{x_{r}y_{r}}\ln|x-y|+S(x,y)\quad\text{with}\,\,S\in W^{1,\infty}_{\rm loc}({\mathbb{R}}^{2}_{+}\times{\mathbb{R}}^{2}_{+}).$
Proof.
We define
$S(x,y):=\mathcal{G}(x,y)-\frac{1}{2\pi}\sqrt{x_{r}y_{r}}\ln|x-y|=-\frac{1}{2\pi}\sqrt{x_{r}y_{r}}\Bigg{(}F\left(\frac{|x-y|^{2}}{x_{r}y_{r}}\right)+\ln|x-y|\Bigg{)}.$
Let $K$ a compact subset of +${\mathbb{R}}^{2}_{+}$, then there exists $R_{0}>0$ such that for any $x,y\in K$ we have $\frac{|x-y|^{2}}{x_{r}y_{r}}\leqslant R_{0}$. Hence, by Lemma 2.4, one has for all $x,y\in K$
$\Bigg{|}F\left(\frac{|x-y|^{2}}{x_{r}y_{r}}\right)+\ln|x-y|\Bigg{|}=\Big{|}\ln(\sqrt{x_{r}y_{r}})+\mathcal{O}(1)\Big{|}\leqslant C_{K}.$
For the derivative, we compute
$\nabla_{x}F\left(\frac{|x-y|^{2}}{x_{r}y_{r}}\right)$ $=\Big{(}\frac{2(x-y)}{x_{r}y_{r}}-\frac{|x-y|^{2}e_{r}}{x_{r}^{2}y_{r}}\Big{)}F^{\prime}\left(\frac{|x-y|^{2}}{x_{r}y_{r}}\right)$
$=-\frac{x-y}{|x-y|^{2}}+\mathcal{O}(1)+\mathcal{O}\left(\frac{1}{x_{r}}\right),$
which implies that
$\Bigg{|}\nabla_{x}F\left(\frac{|x-y|^{2}}{x_{r}y_{r}}\right)+\nabla_{x}\ln|x-y|\Bigg{|}\leqslant C_{K}$
for all $x,y\in K$. As we have the same estimates for $\nabla_{y}$, this ends the proof. ∎
This expansion of the Green’s function together with (2.8) allows one to derive a decomposition of the velocity field $u$. For further reference, we denote the most singular part of the stream function
(2.12) $\psi^{\varepsilon}(t,x)=\psi[\omega^{\varepsilon}](t,x):=\int_{{\mathbb{R}}^{2}_{+}}\frac{\sqrt{x_{r}y_{r}}}{2\pi}\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y.$
As will be detailed later, this quantity plays a central role in determining the energy and the local induction for the propagation of vortex rings.
Lemma 2.6.
Let $K$ a compact subset of +${\mathbb{R}}^{2}_{+}$ and $\omega\in L^{\infty}({\mathbb{R}}_{+}^{2})$ with $\operatorname{supp}(\omega)\subset K$. There exists $C_{K}>0$ depending only on $K$ such that the velocity field $u$ defined by (2.8) satisfies
$u:=u_{K}+u_{L}+u_{R}$
with
(2.13) $\begin{split}u_{K}(x)&=\frac{1}{x_{r}}\int_{{\mathbb{R}}^{2}_{+}}\frac{\sqrt{x_{r}y_{r}}}{2\pi}K(x-y)y_{r}\omega(y)\mathrm{d}y,\quad\text{with }\,K(x,y)=\nabla^{\perp}\ln|x-y|=\frac{(x-y)^{\perp}}{|x-y|^{2}},\\ u_{L}(x)&=-\frac{e_{z}}{4\pi x_{r}}\int_{{\mathbb{R}}^{2}_{+}}\sqrt{\frac{y_{r}}{x_{r}}}\ln|x-y|y_{r}\omega(y)\mathrm{d}y=-\frac{e_{z}}{2x_{r}^{2}}\psi[\omega](x),\\ u_{R}(x)&=\frac{1}{x_{r}}\int_{{\mathbb{R}}^{2}_{+}}\nabla_{x}^{\perp}S(x,y)y_{r}\omega(y)\mathrm{d}y,\end{split}$
where $\|u_{R}\|_{L^{\infty}(K)}\leqslant C_{K}\int y_{r}\omega(y)\mathrm{d}y$.
Proof.
By the decomposition of Corollary 2.5, we have
$\nabla^{\perp}\mathcal{G}(x,y)=\frac{\sqrt{x_{r}y_{r}}}{2\pi}\frac{(x-y)^{\perp}}{|x-y|^{2}}-\frac{1}{4\pi}\sqrt{\frac{{y_{r}}}{{x_{r}}}}\ln|x-y|e_{z}+\nabla^{\perp}_{x}S(x,y),$
hence the decomposition of $u$ comes from (2.8) and (2.10). The desired estimate of $u_{R}$ follows the estimates of $S$ in Corollary 2.5. ∎
Note that the most singular term $u_{K}$ in the decomposition stems from the analogue of the 2D Biot-Savart law. It corresponds to the standard spinning around the filament and does not contribute to the motion of the vortex rings. The second term given by $u_{L}$ is linked to the radial correction of the 2D Biot-Savart law in the present axisymmetric setting. It points in the vertical $e_{z}$-direction and its speed amounts to $\frac{-1}{2x_{r}^{2}}\psi^{\varepsilon}(t,x)$ where the latter is closely linked to local energy, see Lemma 4.6 below. In the absence of strong interactions, it accounts for the leading order displacement of the vortex ring according to the LIA or binormal curvature flow, see (1.4). Finally, the expansion of the Biot-Savart kernel allows one to treat $u_{R}$ as term of lower order.
Remark 2.7.
In dimension reduced settings like the present one, it is well-known that the $3D$-elliptic problem translates to a weighted $2D$ elliptic problem like (2.9). For instance, this also occurs for the (degenerate) lake equations and the $3D$ Euler equations under helicoidal symmetry. A common feature is that the unique solution to the respective weighted elliptic problems is given by
$\Psi(x)=\int_{{\mathbb{R}}^{2}_{+}}\mathcal{G}(x,y)y_{r}\omega(y)\mathrm{d}y$
where the Green function shares similarities with the classical one on ${\mathbb{R}}^{2}$. For the lake equations on bounded domains, it was proven in [19, 2] that
$\mathcal{G}_{\mathrm{lake}}(x,y)=\frac{\sqrt{b(x)b(y)}}{2\pi}\ln|x-y|+S(x,y)\quad\text{with}\quad S\in W^{1,\infty}_{\rm loc}(\Omega^{2}),$
is the Green kernel associated to the following elliptic problem
$\begin{cases}\displaystyle\operatorname{div}\left(\frac{1}{b(x)}\nabla\Psi(x)\right)=b(x)\omega(x)\\ \Psi|_{\partial\Omega}=0.\end{cases}$
Note that the 3D axisymmetric Euler equations without swirl can formally be recovered as special case of the lake equations by setting $b(x_{z},x_{r})=x_{r}$ and considering a lake given by the half-plane. A similar result holds for the $3D$ Euler equations under helicoidal symmetry [20], but with a more complicated expansion due to the fact that the anelastic constraint is not a scalar function $1/b(x)$, but a matrix. A key advantage of such an expansion of the respective Green function lies in the decomposition of the associated velocity field, see Lemma 2.6, that highlights the link to the LIA, namely the contribution of $u_{L}$, see [19, 31, 20].
3. A vortex ring under the influence of the other rings
Our strategy for proving Theorem 1.1 consists in focusing on the dynamics of a single blob of vorticity, by considering the effect of the other blobs as an exterior field. To proceed, we place ourselves in the framework of Theorem 1.1. Specifically, let $X^{*}=(z^{*},r^{*})\in{\mathbb{R}}^{2}_{+}$ and $(Y_{i,0})_{1\leqslant i\leqslant N}$ be $N$ points in ${\mathbb{R}}^{2}$ such that $Y_{i,0,r}\neq Y_{j,0,r}$ for all $i\neq j$ and let $X_{i,0}^{\varepsilon}$ given by (1.6). Finally, let $M_{0}>0$ be fixed and $\omega_{0}^{\varepsilon}\in L^{\infty}({\mathbb{R}}^{2}_{+})$ satisfying (1.7) for every $\varepsilon>0$. Further, we denote by $(u^{\varepsilon},\omega^{\varepsilon})$ the unique weak solution of (1.5) with initial data $\omega_{0}^{\varepsilon}$ on $[0,T_{0})$ where $T_{0}$ is arbitrary large (in the sense of Proposition 2.1). The vorticity can be decomposed as $\omega^{\varepsilon}(t,\cdot)=\sum_{i=1}^{N}\omega_{i}^{\varepsilon}(t,\cdot)$ where $\omega_{i}^{\varepsilon}$ corresponds to the transport of $\omega_{i,0}^{\varepsilon}$, see Remark 2.3. For each $i\in\{1,\ldots,N\}$, we introduce the exterior field $F^{\varepsilon}_{i}:[0,+\infty)\times{\mathbb{R}}^{2}_{+}\to{\mathbb{R}}^{2}$ given by
(3.1) $F_{i}^{\varepsilon}(t,x)=\frac{1}{x_{r}}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\nabla_{x}^{\perp}\mathcal{G}(x,y)y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y,$
so that the $i$-th blob $\omega_{i}^{\varepsilon}$ satisfies the following equations:
(3.2) $\begin{cases}\displaystyle\partial_{t}\,\omega_{i}^{\varepsilon}+\frac{1}{|\ln\varepsilon|}(u_{i}^{\varepsilon}+F_{i}^{\varepsilon})\cdot\nabla\omega_{i}^{\varepsilon}=0&\text{ in }{\mathbb{R}}_{+}^{*}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ u_{i}^{\varepsilon}=\frac{1}{x_{r}}\nabla^{\perp}\Psi_{i}^{\varepsilon}&\text{in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \operatorname{div}\big{(}\frac{1}{x_{r}}\nabla\Psi_{i}^{\varepsilon}\big{)}=x_{r}\omega_{i}^{\varepsilon}&\text{ in }{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}_{+},\quad\Psi_{i}^{\varepsilon}=0\text{ on }{\mathbb{R}}_{+}\times\partial{\mathbb{R}}^{2}_{+},\vspace{1mm}\\ \omega_{i}^{\varepsilon}(0,\cdot)=\omega_{i,0}^{\varepsilon}&\text{ in }{\mathbb{R}}^{2}_{+}.\end{cases}$
As $\operatorname{div}(x_{r}(u_{i}^{\varepsilon}+F_{i}^{\varepsilon}))=0$, it is clear that the total mass of vorticity is conserved:
(3.3) +$\int_{{\mathbb{R}}^{2}_{+}}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}y=\int_{{\mathbb{R}}^{2}_{+}}x_{r}\omega_{i,0}^{\varepsilon}(t,x)\mathrm{d}y=\gamma,\qquad\forall t\in[0,T_{0}].$
For any $r_{0},h_{0},d>0$ and $\varepsilon>0$ we define the rectangles $\mathcal{A}_{\varepsilon}^{r_{0}}(d,h_{0})$ corresponding to hollow cylinders in the 3D setting as
(3.4) $\mathcal{A}_{\varepsilon}^{r_{0}}(d,h_{0})=\left\{(z,r)\in{\mathbb{R}}^{2}_{+},|r-r_{0}|\leqslant\frac{d}{\sqrt{|\ln\varepsilon}|}\text{ and }|z-z^{*}|\leqslant h_{0}\right\}.$
The height $h_{0}>0$ will be chosen later, and we set
(3.5) $d_{0}:=\frac{1}{4}\min_{j\neq k}|Y_{j,0,r}-Y_{k,0,r}|.$
From the assumption (1.7) on $\omega_{0}^{\varepsilon}$, we have for every $i\in\{1,\ldots,N\}$,
$\operatorname{supp}\omega_{i,0}^{\varepsilon}\subset B(X^{\varepsilon}_{i,0},\varepsilon)\subset\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(\frac{d_{0}}{4},\frac{h_{0}}{4}\right)$
as soon as $\varepsilon\leqslant\varepsilon_{0}$ for $\varepsilon_{0}>0$ depending on $h_{0}$.
Let $T_{0}>0$ to be chosen later. Defining
(3.6) $T_{\varepsilon}:=\sup\Big{\{}t\in[0,T_{0}],\,\forall s\in[0,t],\ \forall i\in\{1,\ldots,N\},\ \operatorname{supp}\omega_{i}^{\varepsilon}(s,\cdot)\subset\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)\Big{\}}$
which verifies, by continuity of the trajectories, $T_{\varepsilon}>0$ for every $\varepsilon>0$. Note that in particular, up to choosing $\varepsilon_{0}$ small enough, we have
(3.7) $\operatorname{supp}\omega_{i}^{\varepsilon}(t,\cdot)\subset\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)\subset[z^{*}-h_{0},z^{*}+h_{0}]\times[r^{*}/2,2r^{*}],\\ \quad\forall t\in[0,T_{\varepsilon}],\ i\in\{1,\ldots,N\},\ \varepsilon\in(0,\varepsilon_{0}].$
Next, we prove that the influence on $\omega_{i}^{\varepsilon}$ of the other vortex rings is not too singular. More precisely we have the following lemma.
Lemma 3.1.
Provided $h_{0}\leqslant 1$, there exists $\varepsilon_{0}>0$ depending only on $X^{*}$, $(Y_{i,0})$ and $M_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, for every $i\in\{1,\ldots,N\}$, every $x,y\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)$ and every $t\in[0,T_{\varepsilon}]$,
[…]
with
$C_{F}=\frac{\gamma C^{*}}{d_{0}}\qquad\text{ and }\qquad C_{F}^{\prime}=\frac{\gamma C^{*}}{d_{0}^{2}},$
and where $C^{*}$ depends only on $X^{*}$.
Proof.
First, we note that $\operatorname{div}(x_{r}F_{i}^{\varepsilon}(t,x))=0$ by definition of $F_{i}^{\varepsilon}$ (3.1).
By (3.7), we infer the existence of $R_{0}>0$ such that $\frac{|x-y|^{2}}{x_{r}y_{r}}\leqslant R_{0}$ for all $x,y\in[z^{*}-1,z^{*}+1]\times[r^{*}/2,2r^{*}]$. Thanks to the expression of $\mathcal{G}$ in terms of $F$ (2.11), we deduce from Lemma 2.4 that
$|\nabla\mathcal{G}(x,y)|\leqslant\frac{C}{|x-y|},\quad|\nabla^{2}\mathcal{G}(x,y)|\leqslant\frac{C}{|x-y|^{2}},$
for some $C>0$ depending only on $X^{*}$.
Next we notice that for all $i\neq j$ and $(x,y)\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)\times\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{j,0,r}}\left(d_{0},h_{0}\right)$, we have
$|x-y|\geqslant|x_{r}-y_{r}|\geqslant|X^{\varepsilon}_{i,0,r}-X^{\varepsilon}_{j,0,r}|-|x_{r}-X^{\varepsilon}_{i,0,r}|-|y_{r}-X^{\varepsilon}_{j,0,r}|\geqslant\frac{4d_{0}}{\sqrt{|\ln\varepsilon|}}-\frac{2d_{0}}{\sqrt{|\ln\varepsilon|}}.$
By the definition of $F_{i}$ (3.1) and the conservation of the mass (3.3), this ends the proof. ∎
The remainder of the paper is then devoted to proving that a blob of vorticity $\omega_{i}^{\varepsilon}$ which is transported by $u_{i}^{\varepsilon}$, the vector field generated by $\omega_{i}^{\varepsilon}$ through the Biot-Savart law, see Lemma 2.6, and by an exterior field $F_{i}^{\varepsilon}$, which satisfies Lemma 3.1, remains concentrated in the sense of Theorem 1.1 on $[0,T_{\varepsilon}]$. More precisely, by choosing $T_{0}>0$ and $\varepsilon_{0}>0$ small enough, we will show a stronger localization than the one given in the definition of $T_{\varepsilon}$, i.e.
$\operatorname{supp}\omega_{i}^{\varepsilon}(t,\cdot)\subset\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(\frac{d_{0}}{2},\frac{h_{0}}{2}\right),\qquad\forall t\in[0,T_{\varepsilon}],\ \varepsilon\in(0,\varepsilon_{0}],\ i\in\{1,\ldots,N\}.$
By definition of $T_{\varepsilon}$ (3.6), this will imply that $T_{\varepsilon}=T_{0}$ and that the strong localization remains valid on $[0,T_{0}]$.
In this respect, it is crucial to verify that all the estimates below are uniform with respect to $i$. Since the proofs will focus only on one patch with one exterior field, we will often drop the index $i$ for the sake of readability.
In the sequel, every constant $C,C_{0},\varepsilon_{0}$ may depend on $X^{*}$, $(Y_{i,0})$ and $M_{0}$ and we omit to mention this dependence. At the opposite, we will carefully track the dependence on $h_{0}$ and $T_{0}$. Finally, we consider from now on that $\gamma>0$, and the opposite sign will be reached simply by reversing the time.
4. Weak localization
In this section, we focus on the weak localization which corresponds to Item (ii) in Theorem 1.1. The key proposition here is the analog of [49, Lemma 2.1], also used in a different context in [28].
Proposition 4.1 (Weak concentration).
Provided $T_{0}\leqslant 1/(26C^{\prime}_{F})$ and $h_{0}\leqslant\min(r^{*}/240,1/2)$ the following holds true: for any positive function $\varepsilon\mapsto\rho(\varepsilon)$ such that $\rho(\varepsilon)\sqrt{|\ln\varepsilon|}$ tends to zero as $\varepsilon\to 0$, there exists $\varepsilon_{\rho}>0$, depending only on $\rho$ and $h_{0}$, such that for any $\varepsilon\in(0,\varepsilon_{\rho}]$, every $i\in\{1,\ldots,N\}$ and all $t\in[0,T_{\varepsilon}]$ there is an open set $\Omega_{\rho,i}^{\varepsilon}(t)$ satisfying
$\frac{1}{\gamma}\int_{\Omega_{\rho,i}^{\varepsilon}(t)}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x\geqslant 1-\frac{h_{0}}{r^{*}}\rho(\varepsilon)\ln|\ln\varepsilon|\quad\text{and}\quad\mathrm{diam}\,\Omega_{\rho,i}^{\varepsilon}(t)\leqslant\varepsilon e^{120/\rho(\varepsilon)}.$
We do not claim that the constants are optimal, but we write them here explicitly in order to focus on the dependence with respect to $h_{0}$ and $T_{0}$.
Even if the proof was originally inspired by [49], we need to add several new arguments due to the stronger singularity of our problem. Indeed, the vortex patches are here close to each other, with a distance of order $1/\sqrt{|\ln\varepsilon|}$. The strategy in this section is to first assume that weak localization holds up to some time $\widetilde{T}_{\varepsilon}$, which is less precise than the desired estimates. This will allow to obtain sharp energy estimates and to deduce a sharper weak localization property on $[0,\widetilde{T}_{\varepsilon}]$, which will imply that $\widetilde{T}_{\varepsilon}=T_{\varepsilon}$ and that all the estimates proved up to $\widetilde{T}_{\varepsilon}$ hold actually true up to $T_{\varepsilon}$. Among these estimates and as they will be used later, we state a lemma encompassed in this bootstrap argument, which will be established in the proof of Proposition 4.1. This lemma involves the center of vorticity of each blob, which was defined at relation 1.12.
Lemma 4.2.
Provided $T_{0}\leqslant 1/(26C^{\prime}_{F})$ and $h_{0}\leqslant\min(r^{*}/240,1/2)$, there exists $\varepsilon_{0}$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$,
(4.1) $\frac{1}{\gamma}\int_{{\mathbb{R}}^{2}_{+}}|x-b^{\varepsilon}_{i}(t)|x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x\leqslant 5h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
Finally, at the end of this section, we will deduce from Proposition 4.1 the concentration property around the center of vorticity.
Corollary 4.3.
Provided $T_{0}\leqslant 1/(26C^{\prime}_{F})$ and $h_{0}\leqslant\min(r^{*}/240,1/2)$ the following holds true: there exists $\varepsilon_{0}>0$ depending on $h_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$ and $i\in\{1,\ldots,N\}$,
$\sup_{t\in[0,T_{\varepsilon}]}\left|\gamma-\int_{B\left(b^{\varepsilon}_{i}(t),3h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x\right|\leqslant\gamma\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
All these statements are proved within this section.
4.1. Weak localization bootstrapping time
We begin by introducing a time $\widetilde{T}_{\varepsilon}$ such that the solution is weakly concentrated in a less precise way than in Proposition 4.1. Namely, we set $\widetilde{T}_{\varepsilon}>0$ the largest time $\widetilde{T}_{\varepsilon}\leqslant T_{\varepsilon}$ such that for every $t\in[0,\widetilde{T}_{\varepsilon}]$ and every $i\in\{1,\ldots,N\}$, there exists a domain $\Omega_{i}^{\varepsilon}(t)$ such that
(4.2) $\frac{1}{\gamma}\int_{{\mathbb{R}}^{2}_{+}\setminus\Omega_{i}^{\varepsilon}(t)}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x<\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\quad\text{and}\quad\operatorname{diam}\Omega_{i}^{\varepsilon}(t)<\frac{1}{|\ln\varepsilon|}.$
To check that $\widetilde{T}_{\varepsilon}>0$, we may set $\Omega_{i}^{\varepsilon}(t)=X^{\varepsilon}(t,\operatorname{supp}(\omega_{i,0}^{\varepsilon}))$, so that
+$\int_{{\mathbb{R}}^{2}_{+}\setminus\Omega_{i}^{\varepsilon}(t)}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x=\int_{{\mathbb{R}}^{2}_{+}\setminus\operatorname{supp}(\omega_{i,0}^{\varepsilon})}x_{r}\omega_{i,0}(x)\mathrm{d}x=0,$
and observe that by virtue of (2.6) and (2.7) we have
$\text{diam}(\Omega_{i}^{\varepsilon}(t))\leqslant\text{diam}(\operatorname{supp}(\omega_{i,0}^{\varepsilon}))+C_{\varepsilon}t\leqslant 2\varepsilon+C_{\varepsilon}t,$
with $C_{\varepsilon}$ possibly diverging as $\varepsilon\to 0$, so that $\text{diam}(\Omega_{i}^{\varepsilon}(t))<1/|\ln\varepsilon|$ for $t>0$ sufficiently small.
The different localization properties related to the definition of $T_{\varepsilon}$ (3.6) and $\widetilde{T}_{\varepsilon}$ are illustrated in Figure 1. As often in the rest of this paper, we drop during the proofs the index $i$ in the notation of $\omega_{i}^{\varepsilon}$ and $\Omega_{i}^{\varepsilon}$ since we focus on a single vortex ring in an exterior field.
$r$$z$$X^{\varepsilon}(0)$$X^{\varepsilon}(t)$$2h_{0}$$\frac{2d_{0}}{\sqrt{|\ln\varepsilon|}}$$t\leqslant T_{\varepsilon}$trong localization
[…]
Figure 1. Weak and strong localization: on the left, the support of $\omega^{\varepsilon}$ remains within a disk of size $\mathcal{O}(1)$ for $t\leqslant T_{0}$, and within a strip of width $2d_{0}/\sqrt{|\ln\varepsilon}$ for $t\leqslant T_{\varepsilon}$. On the right, most of the mass is within a circle of diameter $1/|\ln\varepsilon|$ for $t\leqslant\widetilde{T}_{\varepsilon}$, while initially all the mass was within a circle of radius $\varepsilon$.
The first consequence of this definition of $\widetilde{T}_{\varepsilon}$ is that Lemma 4.2 holds true up to $\widetilde{T}_{\varepsilon}$, namely we have the following:
Lemma 4.4.
Provided $h_{0}\leqslant 1$, there exists $\varepsilon_{0}>0$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$, relation (4.1) holds true on the time interval $[0,\widetilde{T}_{\varepsilon}]$. In addition, we have on the same time interval that
$\Omega_{i}^{\varepsilon}(t)\subset B\left(b_{i}^{\varepsilon}(t),\frac{3h_{0}\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right).$
where $\Omega_{i}^{\varepsilon}(t)$ is defined in (4.2).
Proof.
Let $x_{0}\in\Omega^{\varepsilon}(t)\cap\operatorname{supp}(\omega^{\varepsilon}(t,\cdot))$, hence $\Omega^{\varepsilon}(t)\subset B\Big{(}x_{0},\frac{1}{|\ln\varepsilon|}\Big{)}$. By definition of $b^{\varepsilon}$ (1.12) and by (3.3), we write
$\gamma|b^{\varepsilon}(t)-x_{0}|=\left|\int_{{\mathbb{R}}^{2}_{+}}(x-x_{0})x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x\right|\leqslant\int_{{\mathbb{R}}^{2}_{+}}|x-x_{0}|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x.$
We split the integration domain, then using that $\operatorname{diam}(\operatorname{supp}\omega^{\varepsilon}(t))\leqslant 2h_{0}$ (see the definition of $T_{\varepsilon}$ (3.6)) and the definition of $\widetilde{T}_{\varepsilon}$ (4.2) we get that for $\varepsilon$ small enough
+$\int_{{\mathbb{R}}^{2}_{+}}|x-x_{0}|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$ $\leqslant\int_{\Omega^{\varepsilon}(t)}|x-x_{0}|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\int_{{\mathbb{R}}^{2}_{+}\setminus\Omega^{\varepsilon}(t)}|x-x_{0}|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\leqslant\frac{\gamma}{|\ln\varepsilon|}+2h_{0}\gamma\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|},$
and thus
+$\int_{{\mathbb{R}}^{2}_{+}}|x-b^{\varepsilon}(t)|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$ $\leqslant\gamma|x_{0}-b^{\varepsilon}(t)|+\int_{{\mathbb{R}}^{2}_{+}}|x-x_{0}|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\leqslant 2\gamma\left(\frac{1}{|\ln\varepsilon|}+2h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right),$
which implies (4.1) by choosing $\varepsilon_{0}$ small enough such that $1\leqslant\frac{h_{0}}{2}\ln|\ln\varepsilon|$.
Another consequence with this choice of $\varepsilon_{0}$ is that for any $x\in\Omega^{\varepsilon}(t)$,
$|x-b^{\varepsilon}(t)|\leqslant|x-x_{0}|+|x_{0}-b^{\varepsilon}(t)|\leqslant\frac{2}{|\ln\varepsilon|}+2h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\leqslant 3h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
Remark 4.5.
A direct consequence of Lemma 4.4 and definition (4.2) of $\widetilde{T}_{\varepsilon}$ is that, for every $t\in[0,\widetilde{T}_{\varepsilon}]$ and every $i\in\{1,\ldots,N\}$, we have
$\int_{|x-b^{\varepsilon}(t)|>\frac{3h_{0}\ln|\ln\varepsilon|}{|\ln\varepsilon|}}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x\leqslant\gamma\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
4.2. Expansion of the energy
Recalling the definition of $\psi_{i}^{\varepsilon}$ given in (2.12), we first bound it by above and from below.
Lemma 4.6.
Provided $h_{0}\leqslant 1/2$, there exists $\varepsilon_{0}>0$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$, $t\in[0,\widetilde{T}_{\varepsilon}]$ and $x\in\operatorname{supp}\omega^{\varepsilon}_{i}(t)$,
$0\leqslant-\psi^{\varepsilon}_{i}(t,x)\leqslant\frac{\gamma}{2\pi}|\ln\varepsilon|\sqrt{x_{r}b_{i,r}^{\varepsilon}(t)}+\frac{2\gamma}{\pi}h_{0}\ln|\ln\varepsilon|.$
In particular, we get that for $\varepsilon$ small enough,
$|\psi_{i}^{\varepsilon}(t,x)|\leqslant\frac{2\gamma}{\pi}|\ln\varepsilon|r^{*}.$
Proof.
Since $t\leqslant T_{\varepsilon}$, then $\operatorname{diam}\big{(}\operatorname{supp}\omega^{\varepsilon}(t)\big{)}\leqslant 2h_{0}$, so provided that $h_{0}<1/2$, then
$\psi^{\varepsilon}(t,x)\leqslant 0\quad\text{if }x\in\operatorname{supp}\omega^{\varepsilon}(t).$
Next, we observe that
$\psi^{\varepsilon}(t,x)$ $=\frac{\sqrt{x_{r}}}{2\pi}\int\sqrt{b^{\varepsilon}_{r}(t)}\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$
$\hskip 28.45274pt+\frac{\sqrt{x_{r}}}{2\pi}\int\left(\sqrt{y_{r}}-\sqrt{b^{\varepsilon}_{r}(t)}\right)\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$
$:=A_{1}+A_{2}.$
To estimate $A_{1}$, we apply the rearrangement of the mass (see Lemma A.1) to $g(s)=-\ln s\mathds{1}_{(0,1)}$ and $M=(2r^{*})M_{0}/\varepsilon^{2}$, to get
(4.3) $-\int\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$ $\leqslant-2\pi M\int_{0}^{\sqrt{\frac{\gamma}{\pi M}}}s\ln s\mathrm{d}s=-2\pi M\Big{(}\frac{1}{2}s^{2}\ln s-\frac{1}{2}s^{2}\Big{)}\Big{|}_{s=\sqrt{\frac{\gamma}{\pi M}}}$
$\leqslant\gamma|\ln\varepsilon|+\mathcal{O}(1)$
and thus
$-A_{1}\leqslant\frac{\gamma}{2\pi}\sqrt{x_{r}b_{r}^{\varepsilon}(t)}|\ln\varepsilon|+\mathcal{O}(1).$
To estimate $A_{2}$, we recall that for $\varepsilon$ small enough, every $x\in\operatorname{supp}\omega^{\varepsilon}(t)$ satisfies $r^{*}/2\leqslant x_{r}\leqslant 2r^{*}$ which implies that $r^{*}/2\leqslant b_{r}^{\varepsilon}(t)\leqslant 2r^{*}$, then
$|A_{2}|=\left|\frac{\sqrt{x_{r}}}{2\pi}\int\frac{y_{r}-b^{\varepsilon}_{r}(t)}{\sqrt{y_{r}}+\sqrt{b^{\varepsilon}_{r}(t)}}\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y\right|\leqslant\frac{\sqrt{2r^{*}}}{2\pi}\int\frac{|y_{r}-b^{\varepsilon}_{r}(t)|}{2\sqrt{r^{*}/2}}\big{|}\ln|x-y|\big{|}y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y.$
We now split the integral
$|A_{2}|\leqslant$ $\frac{1}{2\pi}\int_{|y-b^{\varepsilon}(t)|\leqslant\frac{3h_{0}\ln|\ln\varepsilon|}{|\ln\varepsilon|}}|y_{r}-b^{\varepsilon}_{r}(t)|\ln|x-y||y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$
$+\frac{1}{2\pi}\int_{|y-b^{\varepsilon}(t)|>\frac{3h_{0}\ln|\ln\varepsilon|}{|\ln\varepsilon|}}|y_{r}-b^{\varepsilon}_{r}(t)|\big{|}\ln|x-y|\big{|}y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$
::=:= $A_{21}+A_{22}.$
We simply bound the first term
$A_{21}\leqslant\frac{3h_{0}}{2\pi}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}(\gamma|\ln\varepsilon|+\mathcal{O}(1))=\frac{3h_{0}\gamma}{2\pi}\ln|\ln\varepsilon|+o(\ln|\ln\varepsilon|).$
For the second term, we use again the mass rearrangement lemma but this time on a function with a mass less than $\gamma\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}$ (see Remark 4.5), to get that
$A_{22}$ $\leqslant\frac{1}{2\pi}\frac{2d_{0}}{\sqrt{|\ln\varepsilon|}}\int_{|y-b^{\varepsilon}(t)|>\frac{3h_{0}\ln|\ln\varepsilon|}{|\ln\varepsilon|}}-\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$
$\leqslant\frac{d_{0}}{\pi\sqrt{|\ln\varepsilon|}}\frac{\gamma\ln|\ln\varepsilon|}{|\ln\varepsilon|}(|\ln\varepsilon|+\mathcal{O}(1))=o(\ln|\ln\varepsilon|).$
This means that
$|A_{2}|\leqslant\frac{3h_{0}\gamma}{2\pi}\ln|\ln\varepsilon|+o(\ln|\ln\varepsilon|),$
and thus that
$-\psi^{\varepsilon}_{i}(t,x)\leqslant\frac{\gamma}{2\pi}|\ln\varepsilon|\sqrt{x_{r}b_{i,r}^{\varepsilon}(t)}+\frac{2\gamma}{\pi}h_{0}\ln|\ln\varepsilon|.$
We now define the total energy related to the $i$th vortex as
(4.4) $\mathcal{E}^{\varepsilon}_{i}(t):=-\iint_{{\mathbb{R}}^{2}_{+}\times{\mathbb{R}}^{2}_{+}}\mathcal{G}(x,y)x_{r}\omega^{\varepsilon}_{i}(t,x)y_{r}\omega^{\varepsilon}_{i}(t,y)\mathrm{d}x\mathrm{d}y.$
Due to the initial concentration, namely (1.7), we have the following.
Lemma 4.7.
We have
$\mathcal{E}_{i}^{\varepsilon}(0)=\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{i,r}^{\varepsilon}(0)+\mathcal{O}(1).$
Proof.
Using the decomposition of the Green kernel given in Corollary 2.5, we write
$\mathcal{E}^{\varepsilon}(0)=$ +$-\iint_{{\mathbb{R}}^{2}_{+}\times{\mathbb{R}}^{2}_{+}}\frac{\sqrt{x_{r}y_{r}}}{2\pi}\ln|x-y|x_{r}\omega^{\varepsilon}_{0}(t,x)y_{r}\omega^{\varepsilon}_{0}(t,y)\mathrm{d}x\mathrm{d}y+\mathcal{O}(1)$
$=$ $-\frac{b_{r}^{\varepsilon}(0)+\mathcal{O}(\varepsilon)}{2\pi}\iint_{{\mathbb{R}}^{2}_{+}\times{\mathbb{R}}^{2}_{+}}\ln|x-y|x_{r}\omega^{\varepsilon}_{0}(t,x)y_{r}\omega^{\varepsilon}_{0}(t,y)\mathrm{d}x\mathrm{d}y+\mathcal{O}(1),$
where we have used that $\operatorname{supp}\omega^{\varepsilon}_{0}\subset B(X^{\varepsilon}_{0},\varepsilon)$. Next we use the mass rearrangement as in (4.3) to state that on the one hand we have
+$\iint_{{\mathbb{R}}^{2}_{+}\times{\mathbb{R}}^{2}_{+}}\ln|x-y|x_{r}\omega^{\varepsilon}_{0}(t,x)y_{r}\omega^{\varepsilon}_{0}(t,y)\mathrm{d}x\mathrm{d}y\geqslant-\gamma^{2}|\ln\varepsilon|+\mathcal{O}(1),$
whereas on the other hand, we use that $|x-y|\leqslant 2\varepsilon$ for $x,y\in\operatorname{supp}\omega^{\varepsilon}_{0}$ to get
+$\iint_{{\mathbb{R}}^{2}_{+}\times{\mathbb{R}}^{2}_{+}}\ln|x-y|x_{r}\omega^{\varepsilon}_{0}(t,x)y_{r}\omega^{\varepsilon}_{0}(t,y)\mathrm{d}x\mathrm{d}y\leqslant\ln(2\varepsilon)\gamma^{2}=-\gamma^{2}|\ln\varepsilon|+\mathcal{O}(1)$
which ends the proof. ∎
We aim to precisely express $\mathcal{E}_{i}^{\varepsilon}(t)$. To this end, we compute its time derivative.
Lemma 4.8.
Provided $h_{0}\leqslant 1/2$ and $T_{0}\leqslant 1/C_{F}^{\prime}$, there exists $\varepsilon_{0}>0$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,\widetilde{T}_{\varepsilon}]$, we have that
$\left|\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}^{\varepsilon}_{i}(t)-\frac{\gamma^{2}}{2\pi}F_{i,r}^{\varepsilon}(t,b^{\varepsilon}_{i}(t))\right|\leqslant\frac{21C^{\prime}_{F}\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|,$
where $C^{\prime}_{F}$ is the lipschitz constant of $(F_{i}^{\varepsilon})_{i}$ (see Lemma 3.1).
Proof.
By the same computations as the ones in the proofs of [20, Lemma 4.10] and [31, Proposition 5.6], using the symmetry of $\mathcal{G}$ and the fact that $\omega^{\varepsilon}$ satisfies (3.2) in the sense of Proposition 2.1, we get
$-\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}^{\varepsilon}(t)=\frac{2}{|\ln\varepsilon|}\iint(u^{\varepsilon}+F^{\varepsilon})(t,x)\cdot\nabla_{x}\mathcal{G}(x,y)x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y.$
We now notice that
$u^{\varepsilon}(t,x)\cdot\int\nabla_{x}\mathcal{G}(x,y)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y=u^{\varepsilon}(t,x)\cdot\big{(}-x_{r}u^{\varepsilon}(t,x)\big{)}^{\perp}=0.$
Therefore,
$-\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}^{\varepsilon}(t)$ $=\frac{2}{|\ln\varepsilon|}\iint F^{\varepsilon}(t,x)\cdot\nabla_{x}\mathcal{G}(x,y)x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$=\frac{1}{\pi|\ln\varepsilon|}\iint F^{\varepsilon}(t,x)\cdot\sqrt{x_{r}y_{r}}\frac{x-y}{|x-y|^{2}}x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$\hskip 28.45274pt+\frac{1}{2\pi|\ln\varepsilon|}\iint F^{\varepsilon}(t,x)\cdot e_{r}\sqrt{\frac{y_{r}}{x_{r}}}\ln|x-y|x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$\hskip 56.9055pt+\frac{2}{|\ln\varepsilon|}\iint F^{\varepsilon}(t,x)\cdot\nabla_{x}S(x,y)x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$:=A_{1}+A_{2}+A_{3}.$
Since $\nabla S\in L^{\infty}([z^{\ast}-h_{0},z^{\ast}+h_{0}]\times[r^{\ast}/2,2r^{\ast}])$ by Corollary 2.5, recalling (3.7) and Lemma 3.1 for $F_{\varepsilon}$, we obtain $A_{3}=\mathcal{O}\big{(}\frac{1}{\sqrt{|\ln\varepsilon}|}\big{)}$.
For $A_{1}$ we use a standard symmetrization technique: we have by exchanging $x$ and $y$
$A_{1}$ $=\frac{1}{2\pi|\ln\varepsilon|}\iint\left(F^{\varepsilon}(t,x)-F_{\varepsilon}(t,y)\right)\cdot\sqrt{x_{r}y_{r}}\frac{x-y}{|x-y|^{2}}x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$|A_{1}|$ $\leqslant\frac{C_{F^{\prime}}}{2\pi}\iint\sqrt{x_{r}y_{r}}x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y\leqslant\frac{r^{\ast}C_{F^{\prime}}\gamma^{2}}{\pi}.$
where we have used Lemma 3.1.
We turn to $A_{2}$. We start by a naive estimate: bounding $F^{\varepsilon}(t,x)$ by $C_{F}\sqrt{|\ln\varepsilon|}$, we get that
$|A_{2}|\leqslant\frac{C_{F}}{\sqrt{|\ln\varepsilon|}}\int\frac{1}{x_{r}}|\psi^{\varepsilon}(t,x)|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x\leqslant\frac{2C_{F}}{r^{*}\sqrt{|\ln\varepsilon|}}\int|\psi^{\varepsilon}(t,x)|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
then using Lemma 4.6, we get that
$A_{2}=\mathcal{O}(\sqrt{|\ln\varepsilon|}),$
and thus
(4.5) $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}^{\varepsilon}(t)=\mathcal{O}(\sqrt{|\ln\varepsilon|}).$
This of course is much less precise than what we are aiming to prove, but we use this information to improve the estimate of $A_{2}$.
Using Lemma 4.7, we get from (4.5) that for every $t\leqslant T_{\varepsilon}\leqslant T_{0}\leqslant 1/C_{F}^{\prime}$,
(4.6) $\mathcal{E}(t)=\frac{\gamma^{2}b_{r}^{\varepsilon}(0)}{2\pi}|\ln\varepsilon|+\mathcal{O}(\sqrt{|\ln\varepsilon|}).$
We now express $A_{2}$ as
$A_{2}=$ $\frac{1}{2\pi|\ln\varepsilon|}F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))\iint\sqrt{\frac{y_{r}}{x_{r}}}\ln|x-y|x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$+\frac{1}{2\pi|\ln\varepsilon|}\iint\big{(}F_{r}^{\varepsilon}(t,x)-F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))\big{)}\sqrt{\frac{y_{r}}{x_{r}}}\ln|x-y|x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$=$ $\frac{1}{|\ln\varepsilon|b_{r}^{\varepsilon}(t)}F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))(-\mathcal{E}^{\varepsilon}(t)+\mathcal{O}(1))$
$+\frac{1}{2\pi|\ln\varepsilon|}F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))\iint\left(\frac{1}{x_{r}}-\frac{1}{b_{r}^{\varepsilon}(t)}\right)\sqrt{x_{r}y_{r}}\ln|x-y|x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
$+\frac{1}{2\pi|\ln\varepsilon|}\iint\big{(}F_{r}^{\varepsilon}(t,x)-F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))\big{)}\sqrt{\frac{y_{r}}{x_{r}}}\ln|x-y|x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
::=:= $A_{21}+A_{22}+A_{23}.$
Using successively the Lipschitz bound on $F^{\varepsilon}$, the fact that $x_{r}>r^{*}/2$, Lemma 4.6 then Lemma 4.4 we get that
$|A_{23}|$ $\leqslant\frac{2C^{\prime}_{F}}{r^{*}}\int|x-b^{\varepsilon}(t)||\psi^{\varepsilon}(t,x)|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\leqslant\frac{2C^{\prime}_{F}}{r^{*}}\frac{2\gamma}{\pi}r^{*}|\ln\varepsilon|\int|x-b^{\varepsilon}(t)|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\leqslant\frac{4C^{\prime}_{F}\gamma^{2}}{\pi}5h_{0}\ln|\ln\varepsilon|.$
On the other hand, using the $L^{\infty}$ bound on $F^{\varepsilon}$, the same arguments as above yield
$|A_{22}|=\mathcal{O}(1).$
We now turn to $A_{21}$. Using the $L^{\infty}$ bound of $F^{\varepsilon}$ and relation (4.6), we get that
$\left|A_{21}+\frac{\gamma^{2}}{2\pi}F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))\right|$ $=\frac{|F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))|}{|\ln\varepsilon|b_{r}^{\varepsilon}(t)}\left|-\mathcal{E}^{\varepsilon}(t)+\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)+\mathcal{O}(1)\right|$
$\leqslant\frac{2C_{F}\gamma^{2}}{r^{*}2\pi\sqrt{|\ln\varepsilon|}}\left(|\ln\varepsilon||b_{r}^{\varepsilon}(0)-b_{r}^{\varepsilon}(t)|+\mathcal{O}(1)\right)$
$=\mathcal{O}(1),$
where we have used that $b^{\varepsilon}(0),b^{\varepsilon}(t)\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)$ (see (3.4) for the definition of this set), hence $|b_{r}^{\varepsilon}(0)-b_{r}^{\varepsilon}(t)|\leqslant 2d_{0}/\sqrt{|\ln\varepsilon|}$.
In conclusion,
$\left|\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{E}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))\right|\leqslant\frac{20\gamma^{2}}{\pi}C^{\prime}_{F}h_{0}\ln|\ln\varepsilon|+\mathcal{O}(1),$
which gives the desired result. ∎
Then, we obtain the following.
Lemma 4.9.
Provided $h_{0}\leqslant 1/2$, there exists $\varepsilon_{0}>0$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,\widetilde{T}_{\varepsilon}]$, we have that
$\left|\frac{\mathrm{d}}{\mathrm{d}t}b_{i,r}^{\varepsilon}(t)-\frac{1}{|\ln\varepsilon|}F_{i,r}^{\varepsilon}(t,b_{i}^{\varepsilon}(t))\right|\leqslant 6C^{\prime}_{F}h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
Proof.
Using (ii) in Proposition 2.1, we compute
$\frac{\mathrm{d}}{\mathrm{d}t}b_{r}^{\varepsilon}(t)$ $=\frac{e_{r}}{\gamma|\ln\varepsilon|}\cdot\int_{{\mathbb{R}}^{2}_{+}}\big{(}u^{\varepsilon}(t,x)+F^{\varepsilon}(t,x)\big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x.$
Using the decomposition of $u^{\varepsilon}$ (2.13) and noticing that $u_{L}^{\varepsilon}\cdot e_{r}=0$, we get that
$\frac{\mathrm{d}}{\mathrm{d}t}b_{r}^{\varepsilon}(t)=$ $\frac{e_{r}}{\gamma|\ln\varepsilon|}\cdot\int_{{\mathbb{R}}^{2}_{+}}u_{K}^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$+\frac{e_{r}}{\gamma|\ln\varepsilon|}\cdot\int_{{\mathbb{R}}^{2}_{+}}u_{R}^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$+\frac{e_{r}}{\gamma|\ln\varepsilon|}\cdot\int F^{\varepsilon}(t,x)x_{r}\omega(t,x)\mathrm{d}x$
::=:= $A_{1}+A_{2}+A_{3}.$
Since $u_{R}^{\varepsilon}$ is uniformly bounded with respect to $\varepsilon$ by Lemma 2.6, we first observe that $A_{2}=\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$. Next, we compute using the expression of $u_{K}^{\varepsilon}$ (2.13) that
$A_{1}=\frac{e_{r}}{2\pi\gamma|\ln\varepsilon|}\cdot\iint\frac{\sqrt{y_{r}}}{\sqrt{x_{r}}}\frac{(x-y)^{\perp}}{|x-y|^{2}}x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
and symmetrize to get that $A_{1}=\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$.
For the term $A_{3}$ we compute that
$A_{3}=\frac{1}{|\ln\varepsilon|}F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))+\frac{1}{\gamma|\ln\varepsilon|}\int\big{(}F_{r}^{\varepsilon}(t,x)-F^{\varepsilon}_{r}(t,b^{\varepsilon}(t))\big{)}x_{r}\omega(t,x)\mathrm{d}x$
then using the Lipschitz hypothesis on $F^{\varepsilon}$ (see Lemma 3.1) with Lemma 4.4, we get that
$\left|A_{3}-\frac{1}{|\ln\varepsilon|}F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))\right|\leqslant\frac{C^{\prime}_{F}}{\gamma}\int|x-b^{\varepsilon}(t)|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x\leqslant 5C^{\prime}_{F}h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
In conclusion,
$\left|\frac{\mathrm{d}}{\mathrm{d}t}b_{r}^{\varepsilon}(t)-\frac{1}{|\ln\varepsilon|}F_{r}^{\varepsilon}(t,b^{\varepsilon}(t))\right|\leqslant 5C^{\prime}_{F}h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}+\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right),$
which ends the proof. ∎
The conclusion of the section is the following.
Corollary 4.10.
Provided $h_{0}\leqslant 1/2$ and $T_{0}\leqslant 1/C_{F}^{\prime}$, there exists $\varepsilon_{0}>0$ depending on $h_{0}$ such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,\widetilde{T}_{\varepsilon}]$, we have that
$\left|\mathcal{E}_{i}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{i,r}^{\varepsilon}(t)\right|\leqslant\frac{25C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|.$
Proof.
Gathering the results of Lemmas 4.8 and 4.9, we have that for every $t\leqslant\widetilde{T}_{\varepsilon}$,
$\left|\frac{\mathrm{d}}{\mathrm{d}t}\left(\mathcal{E}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b^{\varepsilon}_{r}(t)\right)\right|\leqslant\frac{21C^{\prime}_{F}\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|+\frac{3C^{\prime}_{F}\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|.$
There simply remains to integrate in time this relation using Lemma 4.7 to get that
$\left|\mathcal{E}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\right|\leqslant\mathcal{O}(1)+\frac{24C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|.$
4.3. Turkington’s-type Lemma
In this section we prove that Proposition 4.1 holds for any $t\leqslant\widetilde{T}_{\varepsilon}$. More precisely, we prove the following.
Lemma 4.11.
Let $\varepsilon\mapsto\rho(\varepsilon)$ a positive function such that $\rho(\varepsilon)\sqrt{|\ln\varepsilon|}\to 0$ as $\varepsilon\to 0$. Provided $h_{0}\leqslant 1/2$ and $T_{0}\leqslant 1/C_{F}^{\prime}$, there exist $\varepsilon_{\rho}>0$, depending only on $\rho$ and $h_{0}$, such that for any $\varepsilon\in(0,\varepsilon_{\rho}]$, every $i\in\{1,\ldots,N\}$ and all $t\in[0,\widetilde{T}_{\varepsilon}]$, there is an open set $\widetilde{\Omega}_{\rho,i}^{\varepsilon}(t)$ satisfying
+$\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{\rho,i}^{\varepsilon}(t)}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x\leqslant\gamma\frac{h_{0}}{r^{*}}\rho(\varepsilon)\ln|\ln\varepsilon|\quad\text{and}\quad\mathrm{diam}\,\widetilde{\Omega}_{\rho,i}^{\varepsilon}(t)\leqslant\varepsilon e^{B/\rho(\varepsilon)}.$
with
$B=20(5+26C^{\prime}_{F}T_{0})$
which crucially does not depend on $\rho$.
Proof.
For $A>0$ which will be determined later, let us define the open set
$\widetilde{\Omega}_{\rho}^{\varepsilon}(t):=\Big{\{}x\in\operatorname{supp}\omega^{\varepsilon}(t,\cdot),\ -\psi^{\varepsilon}(t,x)\geqslant\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)-\frac{A}{\rho(\varepsilon)}\Big{\}}.$
We have from Corollary 4.10 that
$-\frac{25C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|\leqslant\mathcal{E}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t).$
As
$\mathcal{E}^{\varepsilon}(t)=-\int_{{\mathbb{R}}^{2}_{+}}\psi^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\mathcal{O}(1),$
we obtain that
00 0 $\leqslant\mathcal{E}^{\varepsilon}(t)-\frac{\gamma^{2}}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)+\frac{25C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|$
+$\leqslant\int_{{\mathbb{R}}^{2}_{+}}\Big{(}-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\frac{25C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|+\mathcal{O}(1),$
hence
+$\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}\psi^{\varepsilon}(t,x)+\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x\\ \leqslant\int_{\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\frac{26C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|.$
Outside $\widetilde{\Omega}_{\rho}^{\varepsilon}(t)$, we have by definition of this set:
$\frac{A}{\rho(\varepsilon)}\leqslant\psi^{\varepsilon}(t,x)+\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t).$
It follows that
$\frac{A}{\rho(\varepsilon)}\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\leqslant$ +$\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}\psi^{\varepsilon}(t,x)+\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\leqslant$ $\int_{\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x+\frac{26C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|$
$\leqslant$ $\int_{\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}\sqrt{x_{r}b_{r}^{\varepsilon}(t)}|\ln\varepsilon|\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$+\frac{\gamma}{2\pi}|\ln\varepsilon|\sqrt{b_{r}^{\varepsilon}(t)}\int_{\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}\Big{(}\sqrt{x_{r}}-\sqrt{b_{r}^{\varepsilon}(t)}\Big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$+\frac{26C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|$
$\leqslant$ $\frac{2\gamma^{2}}{\pi}h_{0}\ln|\ln\varepsilon|+\frac{\gamma}{2\pi}|\ln\varepsilon|\frac{\sqrt{2r^{*}}}{2\sqrt{r^{*}/2}}\int_{\widetilde{\Omega}_{\rho}^{\varepsilon}(t)}|x_{r}-b_{r}^{\varepsilon}(t)|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$+\frac{26C^{\prime}_{F}\gamma^{2}}{\pi}T_{0}h_{0}\ln|\ln\varepsilon|$
$\leqslant$ $\frac{\gamma^{2}}{\pi}h_{0}(5+26C^{\prime}_{F}T_{0})\ln|\ln\varepsilon|,$
where we used Lemma 4.6 and Lemma 4.4. Choosing $A=\frac{1}{\pi}\gamma r^{*}(5+26C^{\prime}_{F}T_{0})$ gives the first estimate in the lemma for any $\varepsilon\in(0,\varepsilon_{0}]$.
Inside $\widetilde{\Omega}_{\rho}^{\varepsilon}(t)$, we have
$0\leqslant$ $-\psi^{\varepsilon}(t,x)-\frac{\gamma}{2\pi}|\ln\varepsilon|b_{r}^{\varepsilon}(t)+\frac{A}{\rho(\varepsilon)}=-\frac{1}{2\pi}\int_{{\mathbb{R}}^{2}_{+}}\big{(}\sqrt{x_{r}y_{r}}\ln|x-y|-(\ln\varepsilon)b_{r}^{\varepsilon}(t)\big{)}y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y+\frac{A}{\rho(\varepsilon)}$
$\leqslant$ $-\frac{b_{r}^{\varepsilon}(t)}{2\pi}\int_{{\mathbb{R}}^{2}_{+}}\ln\left(\frac{|x-y|}{\varepsilon}\right)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y-\frac{1}{2\pi}\int_{{\mathbb{R}}^{2}_{+}}(\sqrt{x_{r}y_{r}}-b_{r}^{\varepsilon}(t))\ln|x-y|y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y+\frac{A}{\rho(\varepsilon)}$
$\leqslant$ $-\frac{b_{r}^{\varepsilon}(t)}{2\pi}\int_{{\mathbb{R}}^{2}_{+}}\ln\left(\frac{|x-y|}{\varepsilon}\right)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y+C\sqrt{|\ln\varepsilon|}+\frac{A}{\rho(\varepsilon)},$
where we have used $\sqrt{x_{r}y_{r}}-b_{r}^{\varepsilon}(t)=\mathcal{O}(1/\sqrt{|\ln\varepsilon|})$ for any $x,y\in\operatorname{supp}\omega^{\varepsilon}(t)$ and (4.3). Therefore
+$\int_{{\mathbb{R}}^{2}_{+}\setminus B(x,R\varepsilon)}\ln\left(\frac{|x-y|}{\varepsilon}\right)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y\leqslant$ +$-\int_{{\mathbb{R}}^{2}_{+}\cap B(x,R\varepsilon)}\ln\left(\frac{|x-y|}{\varepsilon}\right)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y+C\sqrt{|\ln\varepsilon|}+\frac{2\pi A}{\rho(\varepsilon)b_{r}^{\varepsilon}(t)}$
$\leqslant$ +$-\int_{{\mathbb{R}}^{2}_{+}\cap B(x,\varepsilon)}\ln\left(\frac{|x-y|}{\varepsilon}\right)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y+C\sqrt{|\ln\varepsilon|}+\frac{4\pi A}{\rho(\varepsilon)r^{*}}$
because $-\ln\frac{|x-y|}{\varepsilon}\leqslant 0$ outside $B(x,\varepsilon)$. By the mass rearrangement Lemma A.1 applied to $g(s)=-\ln\frac{s}{\varepsilon}\mathds{1}_{(0,\varepsilon)}$ and $M=2r^{*}M_{0}/\varepsilon^{2}$, we get
+$-\int_{{\mathbb{R}}^{2}_{+}\cap B(x,\varepsilon)}\ln\left(\frac{|x-y|}{\varepsilon}\right)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y$ $\leqslant-2\pi M\int_{0}^{\min(\varepsilon,\sqrt{\frac{\gamma}{\pi M}})}s\ln\frac{s}{\varepsilon}\mathrm{d}s$
$\leqslant-2\pi M\Big{(}\frac{1}{2}s^{2}\ln\frac{s}{\varepsilon}-\frac{1}{4}s^{2}\Big{)}\Big{|}_{s=\min(\varepsilon,\sqrt{\frac{\gamma}{\pi M}})}=\mathcal{O}(1).$
This implies that
$\ln R\int_{{\mathbb{R}}^{2}_{+}\setminus B(x,R\varepsilon)}y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y\leqslant C\sqrt{|\ln\varepsilon|}+\frac{4\pi A}{\rho(\varepsilon)r^{*}}.$
Choosing $B=A\frac{20\pi}{\gamma r^{*}}$ and $R=\frac{1}{2}e^{B/\rho(\varepsilon)}>1$, we have for $\varepsilon$ small enough (depending on $\rho$ and $A$, hence on $T_{0}$) that
+$\int_{{\mathbb{R}}^{2}_{+}\setminus B(x,R\varepsilon)}y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y\leqslant\frac{4\pi}{r^{*}}\frac{C\rho(\varepsilon)\sqrt{|\ln\varepsilon|}+A}{B-\rho(\varepsilon)\ln 2}\leqslant\frac{5\pi}{r^{*}}\frac{A}{B}=\frac{1}{4}\gamma$
because $\rho(\varepsilon)\sqrt{|\ln\varepsilon|}\to 0$ as $\varepsilon\to 0$. As this inequality holds true for any $x\in\widetilde{\Omega}_{\rho}^{\varepsilon}(t)$, it follows that
$\mathrm{diam}\,\widetilde{\Omega}_{\rho}^{\varepsilon}(t)\leqslant 2R\varepsilon=\varepsilon e^{B/\rho(\varepsilon)}.$
Indeed, if we have $x,\widetilde{x}\in\widetilde{\Omega}_{\rho}^{\varepsilon}(t)$ such that $|x-\widetilde{x}|>2R\varepsilon$, it would mean that
$\int_{B(x,R\varepsilon)\cup B(\widetilde{x},R\varepsilon)}y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y=\int_{B(x,R\varepsilon)}y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y+\int_{B(\widetilde{x},R\varepsilon)}y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}y\geqslant 2\frac{3}{4}\gamma$
which is impossible by the assumption $\int y_{r}\omega^{\varepsilon}(y)\mathrm{d}y=\gamma$ and $\omega^{\varepsilon}\geqslant 0$. ∎
4.4. Bootstrap argument and ends of the proof of the main results of this section
In this paragraph we finally establish that $\widetilde{T}_{\varepsilon}=T_{\varepsilon}$, by proving that for every $t\leqslant\widetilde{T}_{\varepsilon}$ and $i\in\{1,\ldots,N\}$, there exists a domain $\widetilde{\Omega}_{i}^{\varepsilon}(t)$ such that
$\operatorname{diam}\widetilde{\Omega}_{i}^{\varepsilon}(t)\leqslant\frac{1}{2|\ln\varepsilon|},$
and
$\frac{1}{\gamma}\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{i}^{\varepsilon}(t)}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x\leqslant\frac{1}{2}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
By continuity of the trajectories, this will prevent the inequality $\widetilde{T}_{\varepsilon}<T_{\varepsilon}$ from being possible.
By Lemma 4.11 applied to $\rho(\varepsilon)=\frac{B}{|\ln\varepsilon|-2\ln|\ln\varepsilon|}$, there exists a set $\widetilde{\Omega}_{i}^{\varepsilon}(t)$ such that
$\mathrm{diam}\,\widetilde{\Omega}_{i}^{\varepsilon}(t)\leqslant\varepsilon e^{B/\rho(\varepsilon)}=\frac{1}{|\ln\varepsilon|^{2}}\leqslant\frac{1}{2|\ln\varepsilon|}$
for $\varepsilon_{0}$ small enough, and
$\frac{1}{\gamma}\int_{{\mathbb{R}}^{2}_{+}\setminus\widetilde{\Omega}_{i}^{\varepsilon}(t)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x\leqslant\frac{h_{0}}{r^{*}}\rho(\varepsilon)\ln|\ln\varepsilon|=\frac{h_{0}}{r^{*}}\frac{B\ln|\ln\varepsilon|}{|\ln\varepsilon|-2\ln|\ln\varepsilon|}\leqslant\frac{1}{2}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|},$
provided that
(4.7) $\frac{Bh_{0}}{r^{*}}=20(5+26C_{F}^{\prime}T_{0})\frac{h_{0}}{r*}\leqslant\frac{1}{2}.$
The above condition is the constraint about $h_{0}$ and $T_{0}$ in the main results stated at the beginning of this section. Assuming for instance $T_{0}\leqslant 1/(26C_{F}^{\prime})$ and $h_{0}\leqslant r^{*}/240$, all the results of this section reads as: there exists $\varepsilon_{0}>0$ depending only on $h_{0}$ such that for any $\varepsilon\in(0,\varepsilon_{0}]$ and every $i\in\{1,\ldots,N\}$, all the inequalities stated hold true for all $t\in[0,T_{\varepsilon}]$. In particular, Proposition 4.1 and respectively Lemma 4.2 come directly from Lemma 4.11 and resp. Lemma 4.4. Finally, Corollary 4.3 is a direct consequence of Lemma 4.4 combined with the definition of $\Omega_{i}^{\varepsilon}$ (4.2).
Remark 4.12.
The constant $h_{0}$ was chosen small enough so that some constants are absorbed, see (4.7), which explained why we have tracked in all this section the dependence on $h_{0}$. The possibility to choose $h_{0}$ small will not be used anymore, therefore, $h_{0}$ is fixed as $h_{0}=\min(r^{*}/240,1/2)$ and we will not specify the dependence of the constants on $h_{0}$ in the sequel.
Remark 4.13.
Before going to the next part of the paper, let us remark that Proposition 4.1 can give us a much sharper concentration result than Corollary 4.3. Indeed, by taking for instance $\rho(\varepsilon)=\frac{120}{\eta|\ln\varepsilon|}$, for any given $\eta>0$, we get the existence of a domain $\hat{\Omega}^{\varepsilon}_{i}(t)$ such that
$\frac{1}{\gamma}\int_{\hat{\Omega}_{i}^{\varepsilon}(t)}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x\geqslant 1-\frac{h_{0}}{r^{*}}\frac{120}{\eta}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\quad\text{and}\quad\mathrm{diam}\,\hat{\Omega}^{\varepsilon}(t)\leqslant\varepsilon^{1-\eta},$
which means that most of the mass can be found in an area of radius $\varepsilon^{1-\eta}$. However, the estimate on the quantity of outside mass being $\mathcal{O}\left(\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)$, with the localization assumptions we can only prove that
$\operatorname{dist}(b^{\varepsilon}(t),\hat{\Omega}^{\varepsilon}(t))=\mathcal{O}\left(\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right),$
and not better. Therefore, the center of mass can be quite far from where the filament is, compared to the size of the core of the filament, due to the possible filamentation of the filament.
5. Equation of the motion
In Lemma 4.9 together with the proof that $\widetilde{T}_{\varepsilon}=T_{\varepsilon}$ (see the end of the previous section), we have obtained the equation of the motion of the radial component of $b^{\varepsilon}_{i}(t)$. In this section, we now obtain the full equation for both components, then make the link with $(X_{i}^{\varepsilon})_{1\leqslant i\leqslant N}$.
5.1. Motion of the center of mass
Keeping for now the notations of a single filament in an exterior field, we compute the following equation for the motion of the center of mass of the vortex ring.
Lemma 5.1.
Provided $T_{0}\leqslant 1/(26C^{\prime}_{F})$, there exists $\varepsilon_{0}>0$ and $C>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
$\left|\frac{\mathrm{d}}{\mathrm{d}t}b_{i}^{\varepsilon}(t)-\frac{1}{|\ln\varepsilon|}F_{i}^{\varepsilon}(t,b_{i}^{\varepsilon}(t))-\gamma\frac{e_{z}}{4\pi b_{i,r}^{\varepsilon}(t)}\right|\leqslant C(1+T_{0})\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
Proof.
Recall that
$\frac{\mathrm{d}}{\mathrm{d}t}b^{\varepsilon}(t)=\frac{1}{\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}\big{(}u^{\varepsilon}(t,x)+F^{\varepsilon}(t,x)\big{)}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x.$
A few computations were already done in Lemma 4.9, and we obtain that
$\frac{\mathrm{d}}{\mathrm{d}t}b^{\varepsilon}(t)$ $=\frac{1}{\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}u_{K}^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\hskip 28.45274pt+\frac{1}{\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}u_{L}^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\hskip 56.9055pt+\frac{1}{\gamma|\ln\varepsilon|}\int u_{R}^{\varepsilon}(t,x)x_{r}\omega(t,x)\mathrm{d}x$
$\hskip 85.35826pt+\frac{1}{\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}F^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$:=A_{1}+A_{2}+A_{3}+A_{4},$
with $A_{1}=\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$ by symmetrizing, and $A_{3}=\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$.
For the term $A_{4}$ we compute that
$A_{4}=\frac{1}{|\ln\varepsilon|}F^{\varepsilon}(t,b^{\varepsilon}(t))+\frac{1}{\gamma|\ln\varepsilon|}\int\big{(}F^{\varepsilon}(t,x)-F(t,b^{\varepsilon}(t))\big{)}x_{r}\omega(t,x)\mathrm{d}x$
then using from Lemma 3.1 that $F$ is Lipschitz and Lemma 4.2
$\left|A_{4}-\frac{1}{|\ln\varepsilon|}F^{\varepsilon}(t,b^{\varepsilon}(t))\right|\leqslant\frac{C^{\prime}_{F}}{\gamma}\int|x-b^{\varepsilon}(t)|x_{r}\omega(t,x)\leqslant 5C^{\prime}_{F}h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
We now turn to $A_{2}$. Using the definition of $u_{L}$ (2.13), we have
$A_{2}$ $=-\frac{e_{z}}{2\gamma|\ln\varepsilon|}\int\frac{1}{x_{r}^{2}}\psi^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$=-\frac{e_{z}}{2\gamma|\ln\varepsilon||b_{r}^{\varepsilon}(t)|^{2}}\int_{{\mathbb{R}}^{2}_{+}}\psi^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\hskip 28.45274pt-\frac{e_{z}}{2\gamma|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}\left(\frac{1}{x_{r}^{2}}-\frac{1}{|b_{r}^{\varepsilon}(t)|^{2}}\right)\psi^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x.$
By Lemma 4.6 and Lemma 4.2, we have
+$\int_{{\mathbb{R}}^{2}_{+}}\left|\frac{1}{x_{r}^{2}}-\frac{1}{|b_{r}^{\varepsilon}(t)|^{2}}\right||\psi^{\varepsilon}(t,x)|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$ $\leqslant C|\ln\varepsilon|\int_{{\mathbb{R}}^{2}_{+}}\big{|}x_{r}-b_{r}^{\varepsilon}(t)|x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$\leqslant C\ln|\ln\varepsilon|.$
Next we use the definition of the energy (4.4), the expression of $\psi^{\varepsilon}$ (2.12), the decomposition of the kernel provided in Corollary 2.5, with the uniform bound of $S$ in $L^{\infty}_{\rm loc}$ and the support of $\omega^{\varepsilon}(t)$ satisfying (3.7), to write
+$-\int_{{\mathbb{R}}^{2}_{+}}\psi^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x=\mathcal{E}^{\varepsilon}(t)+\mathcal{O}(1),$
then by using Corollary 4.10 we obtain that
$\left|A_{2}-\gamma\frac{e_{z}}{4\pi b_{r}^{\varepsilon}(t)}\right|\leqslant C(1+T_{0})\frac{\ln|\ln\varepsilon|}{\ln\varepsilon}.$
Finally, we show that the velocity created by the other vortices $F_{i}^{\varepsilon}$ behaves as the point vortex interaction in the classical point vortex system (see [40, 41]).
Lemma 5.2.
Provided $T_{0}\leqslant 1/(26C^{\prime}_{F})$, there exists $\varepsilon_{0}>0$ and $C>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have
$\left|F_{i}^{\varepsilon}(t,b_{i}^{\varepsilon}(t))-\frac{\gamma}{2\pi}\sum_{j\neq i}\frac{\big{(}b_{i}^{\varepsilon}(t)-b_{j}^{\varepsilon}(t)\big{)}^{\perp}}{\big{|}b_{i}^{\varepsilon}(t)-b_{j}^{\varepsilon}(t)\big{|}^{2}}\right|\leqslant C\ln|\ln\varepsilon|.$
Proof.
Going back to the definition of the $F_{i}^{\varepsilon}$ (3.1) and the decomposition of $\mathcal{G}$, see Corollary 2.5, we have that
$F_{i}^{\varepsilon}(t,b_{i}^{\varepsilon}(t))$ $=\frac{1}{b_{i,r}^{\varepsilon}(t)}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\nabla^{\perp}_{x}\mathcal{G}(b_{i}^{\varepsilon}(t),y)y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y$
$=\frac{1}{2\pi b_{i,r}^{\varepsilon}(t)}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\sqrt{y_{r}b_{i,r}^{\varepsilon}(t)}\frac{\big{(}b_{i}^{\varepsilon}(t)-y\big{)}^{\perp}}{\big{|}b_{i}^{\varepsilon}(t)-y\big{|}^{2}}y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y$
$\hskip 28.45274pt-\frac{e_{z}}{4\pi b_{i,r}^{\varepsilon}(t)}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\sqrt{\frac{y_{r}}{b_{i,r}^{\varepsilon}(t)}}\ln|b_{i}^{\varepsilon}(t)-y|y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y$
$\hskip 56.9055pt+\frac{1}{b_{i,r}^{\varepsilon}(t)}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\nabla^{\perp}_{x}S(b_{i}^{\varepsilon}(t),y)y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y$
$:=A_{1}+A_{2}+A_{3}.$
By the localization of the vorticity supports entailed in the definition of $T_{\varepsilon}$ (3.6), we know that $|b_{i,r}^{\varepsilon}(t)-X_{i,0,r}|\leqslant d_{0}/\sqrt{|\ln\varepsilon|}$. Hence for $j\neq i$ and $y\in\operatorname{supp}\omega_{j}^{\varepsilon}(t)$ we have by definition of $d_{0}$ in (3.5) that
$|b_{i}^{\varepsilon}(t)-y|\geqslant|X_{i,0,r}-X_{j,0,r}|-|X_{i,0,r}-b_{i,r}^{\varepsilon}(t)|-|y-X_{j,0,r}|\geqslant\frac{2d_{0}}{\sqrt{|\ln\varepsilon|}}.$
It follows that $A_{2}=\mathcal{O}(\ln|\ln\varepsilon|)$. On the other hand, we infer from Corollary 2.5 that $A_{3}=\mathcal{O}(1)$.
We now turn to $A_{1}$. By definition of $T_{\varepsilon}$ (3.6), we have that for $y\in\operatorname{supp}\omega_{j}^{\varepsilon}$ and $j\neq i$
$\frac{|\sqrt{b_{i,r}^{\varepsilon}}-\sqrt{y_{r}}|}{|b_{i}^{\varepsilon}(t)-y|}=\frac{|b_{i,r}^{\varepsilon}-y_{r}|}{(\sqrt{b_{i,r}^{\varepsilon}}+\sqrt{y_{r}})|b_{i}^{\varepsilon}(t)-y|}\leqslant\frac{1}{r^{*}}$
hence
$A_{1}=\frac{1}{2\pi}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\frac{\big{(}b_{i}^{\varepsilon}(t)-y\big{)}^{\perp}}{\big{|}b_{i}^{\varepsilon}(t)-y\big{|}^{2}}y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y+\mathcal{O}(1).$
Moreover, since $|b_{i}^{\varepsilon}-b_{j}^{\varepsilon}|\geqslant|X_{i,0,r}-X_{j,0,r}|-2d_{0}/\sqrt{|\ln\varepsilon|}\geqslant 2d_{0}/\sqrt{|\ln\varepsilon|}$, using that $a/|a|^{2}-b/|b|^{2}=(a-b)/(|a||b|)$, we deduce from Lemma 4.2 that
$\left|A_{1}-\frac{1}{2\pi}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\frac{\big{(}b_{i}^{\varepsilon}-b_{j}^{\varepsilon}\big{)}^{\perp}}{\big{|}b_{i}^{\varepsilon}-b_{j}^{\varepsilon}\big{|}^{2}}y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y\right|$ $\leqslant\frac{1}{2\pi}\sum_{j\neq i}\int_{{\mathbb{R}}^{2}_{+}}\frac{|b_{j}^{\varepsilon}-y|}{|b_{i}^{\varepsilon}-b_{j}^{\varepsilon}||b_{i}^{\varepsilon}-y|}y_{r}\omega_{j}^{\varepsilon}(t,y)\mathrm{d}y+\mathcal{O}(1)$
$\leqslant Ch_{0}|\ln\varepsilon|\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\leqslant C\ln|\ln\varepsilon|.$
Gathering all the previous estimates concludes the proof. ∎
5.2. Limit trajectories
We prove in this section that $(b_{i}^{\varepsilon})_{i}$, defined in (1.12), are located close to $(X_{i}^{\varepsilon})_{i}$ and $(\widetilde{X}_{i}^{\varepsilon})_{i}$, defined in (1.8) and resp. in (1.11).
First, we prove that $(X_{i}^{\varepsilon})_{i}$ and $(\widetilde{X}_{i}^{\varepsilon})_{i}$ are located in $\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}(d_{0},h_{0})$, see (3.4) for the definition of this set, and remains close.
Lemma 5.3.
For $h_{0}$ fixed, there exist positive numbers $T_{X}$, $\varepsilon_{0}$ and $C$, such that for every $t\in[0,T_{X}]$, $\varepsilon\in(0,\varepsilon_{0}]$ and $i\in\{1,\ldots,N\}$, we have
$X_{i}^{\varepsilon}(t),\widetilde{X}_{i}^{\varepsilon}(t)\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}(d_{0},h_{0})\quad\text{and}\quad|X_{i}^{\varepsilon}(t)-\widetilde{X}_{i}(t)|\leqslant\frac{C}{|\ln\varepsilon|}.$
Proof.
First, we consider $\widetilde{Y}_{i}$ defined in (1.10). As it is a system independent of $\varepsilon$, it is clear by continuity of the trajectories that there exists $T_{X}\in(0,1]$ such that
$\widetilde{X}_{i}^{\varepsilon}(t)\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\Big{(}\frac{d_{0}}{2},\frac{h_{0}}{2}\Big{)},\quad\forall t\in[0,T_{X}].$
Next, we remark that $Y_{i}^{\varepsilon}$ is given by
$Y_{i}^{\varepsilon}(t)=\sqrt{|\ln\varepsilon|}\left(X_{i}^{\varepsilon}(t)-X^{*}-\frac{\gamma}{4\pi r^{*}}te_{z}\right),$
which satisfies that $Y_{i}^{\varepsilon}(0)=Y_{i,0}$ and
$\frac{\mathrm{d}}{\mathrm{d}t}Y_{i}^{\varepsilon}(t)$ $=\frac{\gamma}{2\pi}\sum_{j\neq i}\frac{(Y_{i}^{\varepsilon}-Y_{j}^{\varepsilon})^{\perp}}{|Y_{i}^{\varepsilon}-Y_{j}^{\varepsilon}|^{2}}+\sqrt{|\ln\varepsilon|}\frac{\gamma}{4\pi r^{*}}\left(\frac{1}{1+\frac{Y_{i,r}^{\varepsilon}}{r^{*}\sqrt{|\ln\varepsilon|}}}-1\right)e_{z}.$
By continuity of the trajectories, we state that the exists $\widehat{T}_{\varepsilon}\in(0,T_{X}]$ such that $X_{i}^{\varepsilon}(t)\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\Big{(}d_{0},h_{0}\Big{)}$ for all $t\in[0,\widehat{T}_{\varepsilon}]$.
Noticing that
$-\frac{\gamma}{4\pi(r^{*})^{2}}\widetilde{Y}_{i,r}=\sqrt{|\ln\varepsilon|}\frac{\gamma}{4\pi r^{*}}\left(\frac{1}{1+\frac{\widetilde{Y}_{i,r}}{r^{*}\sqrt{|\ln\varepsilon|}}}-1\right)+\mathcal{O}\left(\frac{1}{\sqrt{|\ln\varepsilon|}}\right),$
we conclude that for all $t\in[0,\widehat{T}_{\varepsilon}]$
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{Y}^{\varepsilon}=\mathcal{F}_{1}^{\varepsilon}(\mathcal{Y}^{\varepsilon})$
$\left|\frac{\mathrm{d}}{\mathrm{d}t}\widetilde{\mathcal{Y}}-\mathcal{F}_{1}^{\varepsilon}(\widetilde{\mathcal{Y}})\right|=\mathcal{O}\left(\frac{1}{\sqrt{|\ln\varepsilon|}}\right),$
where
$\mathcal{F}^{\varepsilon}_{1}:\begin{cases}({\mathbb{R}}^{2}_{+})^{N}\to({\mathbb{R}}^{2})^{N}\\ \displaystyle\mathcal{Y}=(Y_{1},\ldots,Y_{N})\mapsto\left(\frac{\gamma}{2\pi}\sum_{j\neq i}\frac{(Y_{i}-Y_{j})^{\perp}}{|Y_{i}-Y_{j}|^{2}}+\sqrt{|\ln\varepsilon|}\frac{\gamma}{4\pi r^{*}}\left(\frac{1}{1+\frac{Y_{i,r}}{r^{*}\sqrt{|\ln\varepsilon|}}}-1\right)e_{z}\right)_{1\leqslant i\leqslant N}.\end{cases}$
The map $\mathcal{F}^{\varepsilon}_{1}$ is Lipschitz continuous when $\mathcal{X}^{\varepsilon},\widetilde{\mathcal{X}}^{\varepsilon}\in\prod_{i}\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)$ with constant $K=\mathcal{O}(1)$, so by Gronwall lemma (see for instance the variant in Lemma A.2), we get that for every $t\leqslant T_{\varepsilon}$,
$\left|\mathcal{Y}^{\varepsilon}(t)-\widetilde{\mathcal{Y}}^{\varepsilon}(t)\right|\leqslant\left(\frac{CT_{X}}{\sqrt{|\ln\varepsilon|}}+0\right)e^{KT_{X}},$
for all $t\in[0,\widehat{T}_{\varepsilon}]$.
In conclusion there exists a constant $C$ depending on $(Y_{i,0})_{i}$ but not on $\varepsilon$ such that for any $t\in[0,\widehat{T}_{\varepsilon}]$,
$|Y_{i}^{\varepsilon}(t)-\widetilde{Y}_{i}(t)|\leqslant\frac{C}{\sqrt{|\ln\varepsilon|}},$
and thus
$|X_{i}^{\varepsilon}(t)-\widetilde{X}_{i}(t)|\leqslant\frac{C}{|\ln\varepsilon|}.$
Choosing $\varepsilon_{0}$ small enough, this allows us to state that
$X_{i}^{\varepsilon}(t)\in\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\Big{(}\frac{3d_{0}}{4},\frac{3h_{0}}{4}\Big{)},$
and hence no blow-up happens on $[0,\hat{T}_{\varepsilon}]$, and by continuity of the trajectories, we get that $\widehat{T}_{\varepsilon}\geqslant T_{X}$, which ends the proof. ∎
We are now in position to prove that $b_{i}^{\varepsilon}$ is close to $X_{i}^{\varepsilon}$.
Proposition 5.4.
Provided $T_{0}\leqslant\min(1/(26C^{\prime}_{F}),T_{X})$, there exist $\varepsilon_{0}>0$ and $C>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
$|b_{i}^{\varepsilon}(t)-X_{i}^{\varepsilon}(t)|\leqslant C\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
Proof.
Let $\mathcal{X}^{\varepsilon}(t)=(X_{1}^{\varepsilon}(t),\ldots,X_{N}^{\varepsilon}(t))$, $\mathcal{B}^{\varepsilon}(t)=(b_{1}^{\varepsilon}(t),\ldots,b_{N}^{\varepsilon}(t))$.
Let
$\mathcal{F}^{\varepsilon}_{2}:\begin{cases}({\mathbb{R}}^{2}_{+})^{N}\to({\mathbb{R}}^{2})^{N}\\ \displaystyle\mathcal{X}=(X_{1},\ldots,X_{N})\mapsto\left(\frac{\gamma}{4\pi X_{i,r}}e_{z}+\frac{\gamma}{2\pi|\ln\varepsilon|}\sum_{j\neq i}\frac{\big{(}X_{i}-X_{j}\big{)}^{\perp}}{\big{|}X_{i}-X_{j}\big{|}^{2}}\right)_{1\leqslant i\leqslant N}.\end{cases}$
Then, we have by definition of the $X_{i}^{\varepsilon}$ that
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{X}^{\varepsilon}(t)=\mathcal{F}^{\varepsilon}_{2}(\mathcal{X}^{\varepsilon}(t))$
and by Lemma 5.1 and Lemma 5.2 that
$\left|\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{B}^{\varepsilon}(t)-\mathcal{F}^{\varepsilon}_{2}(\mathcal{B}^{\varepsilon}(t))\right|\leqslant C\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
By Lemma 5.3, the map $\mathcal{F}^{\varepsilon}_{2}$ is Lipschitz continuous on $\prod_{i}\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(d_{0},h_{0}\right)$ with constant $K=\mathcal{O}(1)$, so by Gronwall lemma (see for instance the variant in Lemma A.2), we get that for every $t\leqslant T_{\varepsilon}$,
$\left|\mathcal{B}^{\varepsilon}(t)-\mathcal{X}^{\varepsilon}(t)\right|\leqslant\left(C\,T_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}+\varepsilon\right)e^{KT_{0}},$
which ends the proof. ∎
From this result, we are now in position to prove Theorem 1.1-$(ii)$ on the time interval $t\in[0,T_{\varepsilon}]$.
Corollary 5.5.
Provided $T_{0}\leqslant\min(1/(26C^{\prime}_{F}),T_{X})$, there exist $\varepsilon_{0}>0$ and $C_{0}>0$ independent of $T_{0}$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$, we have that
$\sup_{t\in[0,T_{\varepsilon}]}\left|\gamma-\int_{B\left(X_{i}^{\varepsilon}(t),C_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x\right|\leqslant\gamma\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
Proof.
By Proposition 5.4 we have that
$B\left(b_{i}^{\varepsilon}(t),3h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)\subset B\left(X_{i}^{\varepsilon}(t),(C+3h_{0})\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)\subset B\left(X_{i}^{\varepsilon}(t),\Big{(}C+\frac{3}{2}\Big{)}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right),$
hence
$B\left(X_{i}^{\varepsilon}(t),\Big{(}C+\frac{3}{2}\Big{)}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)^{c}\subset B\left(b_{i}^{\varepsilon}(t),3h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)^{c},$
and
$\int_{B\left(X_{i}^{\varepsilon}(t),(C+\frac{3}{2})\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)^{c}}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x\leqslant\int_{B\left(b_{i}^{\varepsilon}(t),3h_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\right)^{c}}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x.$
So using Corollary 4.3, we have the desired result by taking $C_{0}=C+3/2$. ∎
Remark 5.6.
It is clear from the previous proof that we have the same result replacing $X_{i}^{\varepsilon}$ by $\widetilde{X}_{i}^{\varepsilon}$, which proves the claim made in the Introduction.
6. Strong localization
All the above results are valid up to $T_{\varepsilon}$, defined in (3.6), in particular Part (ii) of Theorem 1.1. The goal of this section is to get Part (ii) of Theorem 1.1, but also to state that $T_{\varepsilon}=T_{0}$, namely we will prove the following.
Proposition 6.1.
For $h_{0}$ fixed, there exists $T_{0}>0$ and $\varepsilon_{0}>0$ such that for every $\varepsilon\in(0,\varepsilon_{0})$, every $t\leqslant T_{\varepsilon}$ and $i\in\{1,\ldots,N\}$, we have that
$\operatorname{supp}\omega_{i}^{\varepsilon}(t,\cdot)\subset\mathcal{A}_{\varepsilon}^{X^{\varepsilon}_{i,0,r}}\left(\frac{d_{0}}{2},\frac{h_{0}}{2}\right)$
and thus, by the continuity of the trajectories of the fluid particles, we conclude that
$T_{\varepsilon}=T_{0}.$
We start by establishing the strong localization in the radial direction, which requires sharp estimates and a two-step iteration process as in [11], then turn to the vertical strong localization by a simpler version of the same method.
6.1. Strong localization in the radial direction
Let $r_{0}^{i,\varepsilon}=X_{i,0,r}^{\varepsilon}$. For any $R>0$, let
$\mathcal{U}^{i,\varepsilon}_{R}=\{(r,z)\in{\mathbb{R}}^{2}_{+},\;|r-r_{0}^{i,\varepsilon}|<R\},$
and
$m_{t}^{i,\varepsilon}(R)=\int_{(\mathcal{U}^{i,\varepsilon}_{R})^{c}}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x$
and let
$\mu_{t}^{i,\varepsilon}(\eta,R)=\int_{{\mathbb{R}}^{2}}\big{(}1-W_{R,\eta}(x_{r}-r_{0}^{i,\varepsilon})\Big{)}x_{r}\omega^{\varepsilon}_{i}(t,x)\mathrm{d}x$
with
$W_{R,\eta}(s)=\begin{cases}1&\text{ if }|s|\leqslant R\\ 0&\text{ if }|s|\geqslant R+\eta\end{cases}$
satisfying that $W_{R,\eta}\in C^{\infty}$ and $\|W_{R,\eta}^{\prime}\|_{L^{\infty}}\leqslant C/\eta$ and $\|W_{R,\eta}^{\prime\prime}\|_{L^{\infty}}\leqslant C/\eta^{2}$.
We first observe that
(6.1) $\mu_{t}^{i,\varepsilon}(R,\eta)\leqslant m_{t}^{i,\varepsilon}(R)\leqslant\mu_{t}^{i,\varepsilon}(R-\eta,\eta).$
Let $t\leqslant T_{\varepsilon}$, $i\in\{1,\ldots,N\}$ and let $x^{i,\varepsilon,0}\in{\mathbb{R}}^{2}_{+}$ such that
$\Big{|}x^{i,\varepsilon,0}_{r}-r_{0}^{i,\varepsilon}\Big{|}=R_{t}^{i,\varepsilon}:=\max\Big{\{}\Big{|}x_{r}-r_{0}^{i,\varepsilon}\Big{|}\,,\,x\in\operatorname{supp}\omega^{\varepsilon}_{i}(t)\Big{\}}.$
Let $s\mapsto x^{i,\varepsilon,t}(s)$ be the Lagrangian trajectory passing by $x^{i,\varepsilon,0}$ at time $t$, i.e. the solution of
[…]
With these notations, we have the following.
Lemma 6.2.
There exists $C>0$ independent of $T_{0}$ such that
$\frac{\mathrm{d}}{\mathrm{d}s}\Big{[}|x^{i,\varepsilon,t}_{r}(s)-r_{0}^{i,\varepsilon}|\Big{]}(t)\leqslant\frac{C}{\sqrt{|\ln\varepsilon|}}+\frac{C}{|\ln\varepsilon|}\left(\frac{1}{R_{t}^{i,\varepsilon}}+\frac{\sqrt{m_{t}^{i,\varepsilon}(R_{t}^{i,\varepsilon}/2)}}{\varepsilon}\right),$
for all $t\in[0,T_{\varepsilon}]$, $\varepsilon>0$ and $i\in\{0,\ldots,N\}$.
Proof.
We compute that
$\frac{\mathrm{d}}{\mathrm{d}s}\Big{[}|x^{\varepsilon,t}_{r}(s)-r^{\varepsilon}_{0}|\Big{]}(t)=\frac{x^{\varepsilon,t}_{r}(t)-r^{\varepsilon}_{0}}{\big{|}x^{\varepsilon,t}_{r}(t)-r^{\varepsilon}_{0}\big{|}}\frac{e_{r}}{|\ln\varepsilon|}\cdot(u^{\varepsilon}+F^{\varepsilon})(t,x^{\varepsilon,t}(t))$
and thus, recalling that $x^{\varepsilon,t}(t)=x^{\varepsilon,0}$,
$\frac{\mathrm{d}}{\mathrm{d}s}\Big{[}|x^{\varepsilon,t}_{r}(s)-r^{\varepsilon}_{0}|\Big{]}(t)$ $\leqslant\frac{1}{|\ln\varepsilon|}\Big{|}u_{K}^{\varepsilon}(t,x^{\varepsilon,0})+u_{R}^{\varepsilon}(t,x^{\varepsilon,0})+F^{\varepsilon}(t,x^{\varepsilon,0})\Big{|}$
$\leqslant\frac{C}{|\ln\varepsilon|}\int_{{\mathbb{R}}^{2}_{+}}\frac{y_{r}\omega^{\varepsilon}(t,y)}{\big{|}x^{\varepsilon,0}-y\big{|}}\mathrm{d}y+\frac{C}{\sqrt{|\ln\varepsilon|}}.$
We now split the integral on $\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}$ and +$\mathcal{{\mathbb{R}}}^{2}_{+}\setminus\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}.$ If $y\in\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}$, then $|x^{\varepsilon,0}-y|\geqslant R_{t}^{\varepsilon}/2$, and thus
$\int_{\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}}\frac{y_{r}\omega^{\varepsilon}(t,y)}{|x^{\varepsilon,0}-y|}\mathrm{d}y\leqslant\frac{C}{R_{t}^{\varepsilon}}.$
On the complement, we use the following property [^2] See for instance [32] when $\Omega={\mathbb{R}}^{2}$ and $f\in L^{1}\cap L^{\infty}({\mathbb{R}}^{2})$, and next apply this for $f\mathds{1}_{\Omega}$.: for every $\Omega\subset{\mathbb{R}}^{2}$ and $f\in L^{1}\cap L^{\infty}(\Omega)$
$\int_{\Omega}\frac{f(y)}{|x^{\varepsilon,0}-y|}\mathrm{d}y\leqslant C\|f\|_{L^{1}(\Omega)}^{\frac{1}{2}}\|f\|_{L^{\infty}(\Omega)}^{\frac{1}{2}},$
where $C$ is independent of $\Omega$, to obtain that
+$\int_{{\mathbb{R}}^{2}_{+}\setminus\mathcal{U}^{\varepsilon}_{R_{t}^{\varepsilon}/2}}\frac{y_{r}\omega^{\varepsilon}(t,y)}{|x^{\varepsilon,0}-y|}\mathrm{d}y\leqslant\frac{C}{\varepsilon}\sqrt{m_{t}^{\varepsilon}(R_{t}^{\varepsilon}/2)}.$
We now work on estimating the quantity $m_{t}^{i,\varepsilon}(R_{t}^{i,\varepsilon}/2)$. To this end we use a double iterative method inspired by [11].
Lemma 6.3.
There exists $C>0$ independent of $T_{0}$ such that, for every $t\leqslant T_{\varepsilon}$, $\varepsilon>0$, $i\in\{1,\ldots,N\}$, $\eta>0$ and $R\in\left[\frac{d_{0}}{128\sqrt{|\ln\varepsilon|}},\frac{d_{0}}{16\sqrt{|\ln\varepsilon|}}\right]$, we have that
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{i,\varepsilon}(R,\eta)\leqslant C\left(\frac{1}{\eta\sqrt{\ln\varepsilon}}+\frac{m_{t}^{i,\varepsilon}(R/2)}{\eta^{2}|\ln\varepsilon|}\right)m_{t}^{i,\varepsilon}(R)$
Proof.
Recalling that $u_{L}^{\varepsilon}\cdot e_{r}=0$, we compute that
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(R,\eta)=\frac{-1}{|\ln\varepsilon|}\int W_{R,\eta}^{\prime}(x_{r}-r^{\varepsilon}_{0})e_{r}\cdot(u_{K}^{\varepsilon}+u_{R}^{\varepsilon}+F^{\varepsilon})(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x=A_{1}+A_{2}$
where
$A_{1}$ $:=\frac{-1}{|\ln\varepsilon|}\int W_{R,\eta}^{\prime}(x_{r}-r^{\varepsilon}_{0})e_{r}\cdot(u_{R}^{\varepsilon}+F^{\varepsilon})x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
$A_{2}$ $:=\frac{-1}{|\ln\varepsilon|}\int W_{R,\eta}^{\prime}(x_{r}-r^{\varepsilon}_{0})e_{r}\cdot u_{K}^{\varepsilon}x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x.$
We start with $A_{1}$. Recalling that $\|u_{R}^{\varepsilon}\|_{L^{\infty}}\leqslant C$ and $\|F^{\varepsilon}\|_{L^{\infty}}\leqslant C\sqrt{|\ln\varepsilon|}$, and that $W_{R,\eta}^{\prime}=0$ out of $[R,R+\eta]$ we have that
$|A_{1}|\leqslant\frac{C}{\eta\sqrt{|\ln\varepsilon|}}m_{t}^{\varepsilon}(R).$
We now compute $A_{2}$. First, we expend $u_{K}^{\varepsilon}$ and get that
$A_{2}=\frac{-1}{|\ln\varepsilon|}\iint W_{R,\eta}^{\prime}(x_{r}-r^{\varepsilon}_{0})e_{r}\cdot\sqrt{\frac{y_{r}}{x_{r}}}K(x-y)x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
then we symmetrize to get that
$A_{2}=\frac{-1}{|\ln\varepsilon|}e_{r}\cdot\iint f(x,y)K(x-y)x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y$
with
$f(x,y)=\frac{1}{2}\Big{(}W_{R,\eta}^{\prime}(x_{r}-r^{\varepsilon}_{0})\sqrt{\frac{y_{r}}{x_{r}}}-W_{R,\eta}^{\prime}(y_{r}-r^{\varepsilon}_{0})\sqrt{\frac{x_{r}}{y_{r}}}\Big{)}.$
For readability, we omit the subscript $\varepsilon$ in the set of type $\mathcal{U}^{\varepsilon}_{R}$. As $f=0$ on $\mathcal{U}_{R}\times\mathcal{U}_{R}$, we observe that for any $\varphi\in L^{\infty}$
+$\iint_{{\mathbb{R}}^{2}_{+}\times{\mathbb{R}}^{2}_{+}}f\varphi$ =+$=\iint_{{\mathbb{R}}^{2}_{+}\times\mathcal{U}_{R/2}}f\varphi+\iint_{{\mathbb{R}}^{2}_{+}\times\mathcal{U}^{c}_{R/2}}f\varphi$
$=\iint_{\mathcal{U}^{c}_{R}\times\mathcal{U}_{R/2}}f\varphi+\iint_{\mathcal{U}^{c}_{R}\times\mathcal{U}^{c}_{R/2}}f\varphi+\iint_{\mathcal{U}_{R}\times\mathcal{U}_{R/2}^{c}}f\varphi$
and
$\iint_{\mathcal{U}_{R}\times\mathcal{U}_{R/2}^{c}}f\varphi+\iint_{\mathcal{U}_{R}^{c}\times\mathcal{U}_{R}^{c}}f\varphi$ $=\iint_{\mathcal{U}_{R}\times\mathcal{U}_{R}^{c}}f\varphi+\iint_{\mathcal{U}_{R}^{c}\times\mathcal{U}_{R}^{c}}f\varphi=\iint_{{\mathbb{R}}^{2}_{+}\times\mathcal{U}_{R}^{c}}f\varphi$
$=\iint_{\mathcal{U}_{R/2}\times\mathcal{U}_{R}^{c}}f\varphi+\iint_{\mathcal{U}_{R/2}^{c}\times\mathcal{U}_{R}^{c}}f\varphi.$
Summing these two lines, we get that for symmetric $f\varphi$, we may split the integral in three as
+$\iint_{{\mathbb{R}}^{2}_{+}\times{\mathbb{R}}^{2}_{+}}f\varphi=2\iint_{\mathcal{U}^{c}_{R}\times\mathcal{U}_{R/2}}f\varphi+2\iint_{\mathcal{U}^{c}_{R}\times\mathcal{U}^{c}_{R/2}}f\varphi-\iint_{\mathcal{U}^{c}_{R}\times\mathcal{U}^{c}_{R}}f\varphi$
and we denote accordingly $A_{2}:=A_{21}+A_{22}+A_{23}$.
When $(x,y)\in\mathcal{U}^{c}_{R}\times\mathcal{U}_{R/2}$, $|x-y|\geqslant R/2\geqslant\frac{d_{0}}{256\sqrt{|\ln\varepsilon|}}$, hence
$|A_{21}|\leqslant\frac{C}{\eta\sqrt{|\ln\varepsilon|}}m_{t}^{\varepsilon}(R).$
For the two other terms, we use that $|f(x,y)|\leqslant\frac{C}{\eta^{2}}|x-y|$ to get that
$|A_{22}+A_{23}|\leqslant\frac{C}{\eta^{2}|\ln\varepsilon|}\iint_{A_{R}^{c}\times A_{R/2}^{c}}x_{r}\omega^{\varepsilon}(t,x)y_{r}\omega^{\varepsilon}(t,y)\mathrm{d}x\mathrm{d}y\leqslant\frac{C}{\eta^{2}|\ln\varepsilon|}m_{t}^{\varepsilon}(R)m_{t}^{\varepsilon}(R/2),$
which ends the proof. ∎
Lemma 6.4 (First iteration).
For every $\ell>0$, there exists $T_{0,\ell}>0$ and $\varepsilon_{\ell}>0$ depending on $\ell$, such that for every $\varepsilon\in(0,\varepsilon_{\ell}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
$m_{t}^{i,\varepsilon}\left(\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}\right)\leqslant\frac{1}{|\ln\varepsilon|^{\ell}}.$
Proof.
As the dynamics of $(\widetilde{Y}_{i})$ is independent of $\varepsilon$ (see (1.10)), it is clear that we can state that
$|\widetilde{X}_{r}^{\varepsilon}(t)-r^{\varepsilon}_{0}|\leqslant\frac{d_{0}}{300\sqrt{|\ln\varepsilon|}},\quad\forall t\in[0,T_{\varepsilon}]$
by choosing $T_{0}$ small enough. From Lemma 5.3 and up to choose $\varepsilon$ small enough, we have
$|X_{r}^{\varepsilon}(t)-r^{\varepsilon}_{0}|\leqslant\frac{d_{0}}{256\sqrt{|\ln\varepsilon|}}\quad\forall t\in[0,T_{\varepsilon}].$
For such $T_{0}$ and $\varepsilon_{0}$, we notice that
$\mathcal{U}_{R/2}^{\varepsilon c}\subset B\Big{(}X^{\varepsilon}(t),C_{0}\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}\Big{)}^{c}$
provided $R\in\left[\frac{d_{0}}{128\sqrt{|\ln\varepsilon|}},\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}\right]$ and $\varepsilon_{0}$ is small enough. Corollary 5.5 then implies that
$m_{t}^{\varepsilon}(R/2)\leqslant\gamma\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|}.$
Plugging this into the estimates of Lemma 6.3, we get that
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(R,\eta)\leqslant C\left(\frac{1}{\eta\sqrt{\ln\varepsilon}}+\frac{\ln|\ln\varepsilon|}{\eta^{2}|\ln\varepsilon|^{2}}\right)m_{t}^{\varepsilon}(R),$
for any $R\in\left[\frac{d_{0}}{128\sqrt{|\ln\varepsilon|}},\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}\right]$. Now we set $R_{0}=\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}$, $n=[\ln|\ln\varepsilon|]-1$ and $\eta=\frac{R_{0}}{2n}$ and we obtain that
$\frac{1}{\eta\sqrt{|\ln\varepsilon|}}+\frac{\ln|\ln\varepsilon|}{\eta^{2}|\ln\varepsilon|^{2}}=\frac{2n}{R_{0}\sqrt{|\ln\varepsilon|}}+\frac{4n^{2}\ln|\ln\varepsilon|}{R^{2}_{0}|\ln\varepsilon|^{2}}\leqslant Cn+Cn\frac{(\ln|\ln\varepsilon|)^{2}}{|\ln\varepsilon|}\leqslant Cn.$
In conclusion,
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(R,\eta)\leqslant C\,n\,m_{t}^{\varepsilon}(R).$
Integrating, we get that
$\mu_{t}^{\varepsilon}(R,\eta)\leqslant\mu_{0}(R,\eta)+C\,n\int_{0}^{t}m_{s}(R)\mathrm{d}s$
and using (6.1) we get that
$\mu_{t}^{\varepsilon}(R,\eta)\leqslant\mu_{0}(R,\eta)+Cn\int_{0}^{t}\mu_{s}(R-\eta,\eta)\mathrm{d}s,$
provided $R\in\left[\frac{R_{0}}{2},R_{0}\right]$.
Applying this inequality twice we get that
$\mu_{t}^{\varepsilon}(R_{0},\eta)\leqslant\mu_{0}^{\varepsilon}(R_{0},\eta)+Cnt\mu_{0}^{\varepsilon}(R_{0}-\eta,\eta)+(Cn)^{2}\int_{0}^{t}\int_{0}^{s_{1}}\mu_{s_{2}}^{\varepsilon}(R_{0}-2\eta,\eta)\mathrm{d}s_{2}\mathrm{d}s_{1}$
and iterating $n$ times eventually leads to
$\mu_{t}^{\varepsilon}(R_{0},\eta)\leqslant\sum_{j=0}^{n-1}\mu_{0}^{\varepsilon}(R_{0}-j\eta,\eta)(Cnt)^{j}+(Cn)^{n}\int_{0}^{t}\ldots\int_{0}^{s_{n-1}}\mu_{s_{n}}^{\varepsilon}(R_{0}/2,\eta)\mathrm{d}s_{n}\ldots\mathrm{d}s_{1}$
and therefore, since $\mu_{0}(0/2)=0$ and $\mu_{s}^{\varepsilon}\leqslant\gamma$ for every $s$, we obtain that
$\mu_{t}^{\varepsilon}(R_{0},\eta)\leqslant\frac{(Cnt)^{n}}{n!}.$
Using Stirling’s formula we have that for any $\ell>0$,
$|\ln\varepsilon|^{\ell}\mu_{t}^{\varepsilon}(R_{0},\eta)$ $\leqslant\frac{1}{\sqrt{\pi n}}\exp\Big{(}n\big{(}\ln(Cnt)-\ln(n)+1\big{)}+\ell\ln|\ln\varepsilon|\Big{)}$
$\leqslant\frac{2}{\sqrt{\pi\ln|\ln\varepsilon|}}\exp\Big{(}\ln|\ln\varepsilon|\big{(}\ln(Ct)+1+\ell\big{)}\Big{)}.$
We see that if $T_{0,\ell}\leqslant\frac{1}{Ce^{\ell+1}}$, then this expression goes to 0 as $\varepsilon\to 0$ uniformly in $t\in[0,T_{\varepsilon}]$. ∎
Lemma 6.5 (Second iteration).
For every $\ell>0$, there exists $T_{0,\ell}>0$ and $\varepsilon_{\ell}>0$ depending on $\ell$, such that for every $\varepsilon\in(0,\varepsilon_{\ell}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
$m_{t}^{i,\varepsilon}\left(\frac{d_{0}}{16\sqrt{|\ln\varepsilon|}}\right)\leqslant\varepsilon^{\ell}.$
Proof.
Applying Lemma 6.4 for $\ell=1$, we state that for every $R\in\left[\frac{d_{0}}{32\sqrt{|\ln\varepsilon|}},\frac{d_{0}}{16\sqrt{|\ln\varepsilon|}}\right]$,
$m_{t}^{\varepsilon}(R/2)\leqslant m_{t}^{\varepsilon}\Big{(}\frac{d_{0}}{64\sqrt{|\ln\varepsilon|}}\Big{)}\leqslant\frac{1}{|\ln\varepsilon|},$
so by injecting this into Lemma 6.3, we get
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(R,\eta)\leqslant C\left(\frac{1}{\eta\sqrt{|\ln\varepsilon|}}+\frac{1}{\eta^{2}|\ln\varepsilon|^{2}}\right)m_{t}^{\varepsilon}(R).$
By taking this time $R_{0}=\frac{d_{0}}{16\sqrt{|\ln\varepsilon|}}$, $\eta=\frac{R}{2n}$ and $n=[|\ln\varepsilon|]-1$, we get again that
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(R,\eta)\leqslant Cnm_{t}^{\varepsilon}(R).$
Then by the same iterative method than in Lemma 6.4, we obtain that
$\mu_{t}^{\varepsilon}(R_{0},\eta)\leqslant\frac{(Cnt)^{n}}{n!}.$
We then compute that
$\varepsilon^{-\ell}\mu_{t}^{\varepsilon}(R_{0},\eta)$ $\leqslant\frac{1}{\sqrt{\pi n}}\exp\Big{(}n\big{(}\ln(Cnt)-\ln(n)+1\big{)}+\ell|\ln\varepsilon|\Big{)}$
$\leqslant\frac{2}{\sqrt{\pi|\ln\varepsilon|}}\exp\Big{(}|\ln\varepsilon|\big{(}\ln(Ct)+1+\ell\big{)}\Big{)}.$
which again goes to 0 if $T_{0,\ell}\leqslant\frac{1}{Ce^{\ell+1}}$. ∎
We are now able to conclude the radial part of Proposition 6.1.
Lemma 6.6.
There exist $T_{0}>0$ and $\varepsilon_{0}>0$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
$R_{t}^{i,\varepsilon}<\frac{d_{0}}{2\sqrt{|\ln\varepsilon|}}.$
Proof.
Assume that there exist $i\in\{1,\ldots,N\}$ and a time $t_{1}\in(0,T_{\varepsilon}]$ such that $R_{t_{1}}^{\varepsilon}=\frac{d_{0}}{2\sqrt{|\ln\varepsilon|}}$. Then by continuity of $t\mapsto R_{t}^{\varepsilon}$, there exists a time $t_{0}<t_{1}$ such that
$R_{t_{0}}^{\varepsilon}=\frac{d_{0}}{8\sqrt{|\ln\varepsilon|}}\qquad\text{ and}\qquad\forall t\in[t_{0},t_{1}],\,R_{t}^{\varepsilon}\geqslant\frac{d_{0}}{16\sqrt{|\ln\varepsilon|}}.$
Let $C_{1}$ be the constant in Lemma 6.2. Let $f:[t_{0},t_{1}]\to{\mathbb{R}}$ be the solution of
$\begin{cases}f(t_{0})=\frac{d_{0}}{4\sqrt{|\ln\varepsilon|}}\vspace{1mm}\\ \displaystyle f^{\prime}(t)=2\frac{C_{1}}{\sqrt{|\ln\varepsilon|}}+2\frac{C_{1}}{|\ln\varepsilon|}\left(\frac{1}{f(t)}+\frac{\sqrt{m_{t}^{\varepsilon}(R_{t}^{\varepsilon}/2)}}{\varepsilon}\right).\end{cases}$
Then for every $t\in[t_{0},t_{1}]$,
$R_{t}^{\varepsilon}<f(t).$
Indeed, if it isn’t the case, by letting $t_{2}\in[t_{0},t_{1}]$ be the first time when $R_{t_{2}}^{\varepsilon}=f(t_{2})$, then there exists a point $x^{\varepsilon,0}\in\operatorname{supp}\omega^{\varepsilon}(t_{2})$ such that $|x^{\varepsilon,0}-r^{\varepsilon}_{0}|=R_{t_{2}}^{\varepsilon}$. However Lemma 6.2 is then in contradiction with the fact that $t_{2}$ is the first such time, due to the fact that $f(t_{0})>R_{t_{0}}^{\varepsilon}$ and
$f^{\prime}(t_{2})>\frac{C_{1}}{\sqrt{|\ln\varepsilon|}}+\frac{C_{1}}{|\ln\varepsilon|}\left(\frac{1}{R_{t_{2}}^{\varepsilon}}+\frac{\sqrt{m_{t}^{\varepsilon}(R_{t_{2}}^{\varepsilon}/2)}}{\varepsilon}\right).$
Since for every $t\in[t_{0},t_{1}]$, $f(t)>R_{t}^{\varepsilon}\geqslant\frac{d_{0}}{16\sqrt{|\ln\varepsilon|}}$, we can apply Lemma 6.5 with $\ell=2$ to get that for $\varepsilon>0$ small enough,
$f^{\prime}(t)\leqslant\frac{C}{\sqrt{|\ln\varepsilon|}}.$
Therefore, by integrating in time,
$R_{t_{1}}^{\varepsilon}<f(t_{1})\leqslant f(t_{0})+(t_{1}-t_{0})\frac{C}{\sqrt{|\ln\varepsilon|}}\leqslant\frac{d_{0}}{2\sqrt{|\ln\varepsilon|}}$
provided $T_{0}$ is small enough. This is in contradiction with the fact that $R_{t_{1}}^{\varepsilon}=\frac{d_{0}}{2\sqrt{|\ln\varepsilon|}}$. By continuity of the trajectories, the conclusion of the lemma follows from the non-existence of such time $t_{1}$. ∎
6.2. Strong localization in the vertical direction
We now perform a very similar analysis, but in the direction $z$. We introduce this time
$\mathcal{U}_{Z}^{i,\varepsilon}=\{(r,z)\in{\mathbb{R}}^{2}_{+},\;|z-z^{*}|<Z\}$
and
$m_{t}^{i,\varepsilon}(Z)=\int_{(\mathcal{U}_{Z}^{i,\varepsilon})^{c}}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x$
and let
$\mu_{t}^{i,\varepsilon}(\eta,Z)=\int_{{\mathbb{R}}^{2}}\big{(}1-W_{Z,\eta}(x_{z}-z^{*})\Big{)}x_{r}\omega_{i}^{\varepsilon}(t,x)\mathrm{d}x.$
Let $t\leqslant T_{\varepsilon}$, $i\in\{1,\ldots,N\}$ and let $x^{i,\varepsilon,0}\in{\mathbb{R}}^{2}_{+}$ such that
$\Big{|}x^{i,\varepsilon,0}_{z}-z^{*}\Big{|}=Z_{t}^{i,\varepsilon}:=\max\Big{\{}\Big{|}x_{z}-z^{*}\Big{|}\,,\,x\in\operatorname{supp}\omega_{i}^{\varepsilon}(t)\Big{\}}.$
We then have the following.
Lemma 6.7.
There exists $C>0$ independent of $T_{0}$ such that
$\frac{\mathrm{d}}{\mathrm{d}s}\Big{[}|x^{\varepsilon,t}_{z}(s)-z^{*}|\Big{]}(t)\leqslant C+\frac{C}{|\ln\varepsilon|}\left(\frac{1}{Z_{t}^{i,\varepsilon}}+\frac{\sqrt{m_{t}^{i,\varepsilon}(Z_{t}^{i,\varepsilon}/2)}}{\varepsilon}\right),$
for all $t\in[0,T_{\varepsilon}]$, $\varepsilon>0$ and $i\in\{0,\ldots,N\}$.
Proof.
We proceed similarly than in Lemma 6.2, the main difference being that the term $u_{L}^{\varepsilon}\cdot e_{z}$ does not vanish. We compute that
$\frac{\mathrm{d}}{\mathrm{d}s}\Big{[}|x^{\varepsilon,t}_{z}(s)-z^{*}|\Big{]}(t)=\frac{x^{\varepsilon,t}_{z}(t)-z^{*}}{\big{|}x^{\varepsilon,t}_{z}(t)-z^{*}\big{|}}\frac{e_{z}}{|\ln\varepsilon|}\cdot(u^{\varepsilon}+F^{\varepsilon})(t,x^{\varepsilon,t}(t))$
and thus
$\frac{\mathrm{d}}{\mathrm{d}s}\Big{[}|x^{\varepsilon,t}_{z}(s)-z^{*}|\Big{]}(t)\leqslant\frac{1}{|\ln\varepsilon|}\Big{|}u_{K}^{\varepsilon}(t,x^{\varepsilon,0})+u_{L}^{\varepsilon}(t,x^{\varepsilon,0})+u_{R}^{\varepsilon}(t,x^{\varepsilon,0})+F^{\varepsilon}(t,x^{\varepsilon,0})\Big{|}.$
We already know from the proof of Lemma 6.2 that
$\frac{1}{|\ln\varepsilon|}\Big{|}u_{K}^{\varepsilon}(t,x^{\varepsilon,0})+u_{R}^{\varepsilon}(t,x^{\varepsilon,0})+F^{\varepsilon}(t,x^{\varepsilon,0})\Big{|}\leqslant\frac{C}{\sqrt{|\ln\varepsilon|}}+\frac{C}{|\ln\varepsilon|}\left(\frac{1}{Z_{t}^{\varepsilon}}+\frac{\sqrt{m_{t}^{\varepsilon}(Z_{t}^{\varepsilon}/2)}}{\varepsilon}\right).$
We now use the relation (2.13) between $u_{L}^{\varepsilon}$ and $\psi^{\varepsilon}$, the estimate of $\psi^{\varepsilon}$ (see Lemma 4.6) and the fact that $\widetilde{T}_{\varepsilon}=T_{\varepsilon}$ (see Section 4.4) to state that
$|u_{L}^{\varepsilon}(t,x^{\varepsilon,0})|\leqslant C|\ln\varepsilon|$
which allows us to conclude. ∎
Lemma 6.8.
There exists $C>0$ independent of $T_{0}$ such that, for every $t\leqslant T_{\varepsilon}$, $\varepsilon>0$, $i\in\{1,\ldots,N\}$, $\eta>0$ and $Z\in\left[\frac{h_{0}}{128},\frac{h_{0}}{16}\right]$, we have that
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{i,\varepsilon}(Z,\eta)\leqslant C\left(\frac{1}{\eta}+\frac{m_{t}^{i,\varepsilon}(Z/2)}{\eta^{2}|\ln\varepsilon|}\right)m_{t}^{i,\varepsilon}(Z)$
Proof.
We compute that
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}(Z,\eta)=\frac{-1}{|\ln\varepsilon|}\int W_{Z,\eta}^{\prime}(x_{z}-z^{*})e_{z}\cdot(u_{K}^{\varepsilon}+u_{L}^{\varepsilon}+u_{R}^{\varepsilon}+F^{\varepsilon})(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x=A_{1}+A_{2}+A_{3}$
where
$|A_{1}|+|A_{2}|\leqslant C\left(\frac{1}{\eta\sqrt{|\ln\varepsilon|}}+\frac{m_{t}^{\varepsilon}(Z/2)}{\eta^{2}|\ln\varepsilon|}\right)m_{t}^{\varepsilon}(Z)$
by reproducing the arguments of Lemma 6.3 with the only difference that $Z/2\geqslant\frac{h_{0}}{256}$, and
$A_{3}:=\frac{-1}{|\ln\varepsilon|}\int W_{Z,\eta}^{\prime}(x_{z}-z^{*})e_{z}\cdot u_{L}^{\varepsilon}(t,x)x_{r}\omega^{\varepsilon}(t,x)\mathrm{d}x$
satisfies
$|A_{3}|\leqslant C/\eta.$
We only need one iteration process to conclude in that case.
Lemma 6.9.
For every $\ell>0$, there exists $T_{0,\ell}>0$ and $\varepsilon_{\ell}>0$ depending on $\ell$, such that for every $\varepsilon\in(0,\varepsilon_{\ell}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
$m_{t}^{i,\varepsilon}\left(\frac{h_{0}}{16}\right)\leqslant\varepsilon^{\ell}.$
Proof.
As in the beginning of the proof of Lemma 6.4, we remark that, up to choose $T_{0}$ and $\varepsilon_{0}$ small enough, Corollary 5.5 implies that
$m_{t}^{\varepsilon}(Z/2)\leqslant\gamma\frac{\ln|\ln\varepsilon|}{|\ln\varepsilon|},$
provided $Z\in\left[\frac{h_{0}}{64},\frac{h_{0}}{16}\right]$. Injecting this into Lemma 6.8, we get that
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(Z,\eta)\leqslant C\left(\frac{1}{\eta}+\frac{\ln|\ln\varepsilon|}{\eta^{2}|\ln\varepsilon|^{2}}\right)m_{t}^{\varepsilon}(Z).$
We take now $Z_{0}=\frac{d_{0}}{16}$ $\eta=\frac{Z_{0}}{2n}$ with $n=[|\ln\varepsilon|]-1$, and get that
$\frac{\mathrm{d}}{\mathrm{d}t}\mu_{t}^{\varepsilon}(Z,\eta)\leqslant Cnm_{t}^{\varepsilon}(Z),$
for all $Z\in\left[\frac{h_{0}}{64},\frac{h_{0}}{16}\right]$. Then by the same iterative method than in Lemma 6.4, we obtain that
$\mu_{t}^{\varepsilon}(Z_{0},\eta)\leqslant\frac{(Cnt)^{n}}{n!}.$
As in Lemma 6.5 we conclude that
$\varepsilon^{-\ell}\mu_{t}^{\varepsilon}(Z_{0},\eta)$ $\leqslant\frac{1}{\sqrt{\pi n}}\exp\Big{(}n\big{(}\ln(Cnt)-\ln(n)+1\big{)}+\ell|\ln\varepsilon|\Big{)}$
$\leqslant\frac{2}{\sqrt{\pi|\ln\varepsilon|}}\exp\Big{(}|\ln\varepsilon|\big{(}\ln(Ct)+1+\ell\big{)}\Big{)}$
which gives the result as soon as $CT_{0,\ell}\leqslant 1/e^{1+\ell}$. ∎
We then conclude the vertical part of Proposition 6.1.
Lemma 6.10.
There exist $T_{0}>0$ and $\varepsilon_{0}>0$, such that for every $\varepsilon\in(0,\varepsilon_{0}]$, $i\in\{1,\ldots,N\}$ and $t\in[0,T_{\varepsilon}]$, we have that
$Z_{t}^{i,\varepsilon}<\frac{h_{0}}{2}.$
Proof.
Assume that this is false, we obtain this time that $Z_{t}^{\varepsilon}<f(t)$ for $t\in[t_{0},t_{1}]$ where
$\begin{cases}f(t_{0})=\displaystyle\frac{h_{0}}{4},\vspace{1mm}\\ f^{\prime}(t)=\displaystyle 2C_{2}+\frac{2C_{2}}{|\ln\varepsilon|}\left(\frac{1}{f(t)}+\frac{\sqrt{m_{t}^{\varepsilon}(Z_{t}^{\varepsilon}/2)}}{\varepsilon}\right)\end{cases}$
and $f(t)\geqslant\frac{h_{0}}{16}$, and thus
$f^{\prime}(t)\leqslant C,$
so that
$f(t)\leqslant\frac{h_{0}}{4}+C(t_{1}-t_{0}).$
Taking $T_{0}$ small enough leads to the contradiction and concludes the proof. ∎
6.3. Conclusion of the proof of Theorem 1.1
By Lemmas 6.6 and 6.10, we have obtained the proof of Proposition 6.1 which implies to Part (ii) of Theorem 1.1, see (3.4) for the definition of $\mathcal{A}_{\varepsilon}^{r_{0}}(d,h_{0})$.
Proposition 6.1 also implies that $T_{\varepsilon}=T_{0}$, so all the estimates we obtained throughout this article hold on the time interval $[0,T_{0}]$ which is independent of $\varepsilon$. In particular Corollary 5.5 gives Part (i) of Theorem 1.1.
Acknowledgements. This work was supported by the BOURGEONS project, grant ANR-23-CE40-0014-01 of the French National Research Agency (ANR). L.E.H. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project- ID 258734477 – SFB 1173. C.L. also benefited of the support of the ANR under France 2030 bearing the reference ANR-23-EXMA-004 (Complexflows project of the PEPR MathsViVEs).
Appendix A Two classical lemmas
In this section, we state two classical lemmas which are used in the article.
First, we recall a standard lemma for mass rearrangement, whose proof is available in [31, Lemma B.1]. Let $g$ be a non increasing continuous function from $(0,+\infty)$, non-negative, such that $s\mapsto sg(s)\in L^{1}_{\mathrm{loc}}\big{(}[0,\infty)\big{)}$. Let
[…]
Lemma A.1.
For all $x\in{\mathbb{R}}^{2}$, we have
$\max_{f\in\mathcal{E}_{M,\gamma}}\int_{{\mathbb{R}}^{2}}g\big{(}|x-y|\big{)}f(y)\mathrm{d}y=2\pi M\int_{0}^{R}sg(s)\mathrm{d}s$
where $R=\sqrt{\frac{\gamma}{\pi M}}$, namely that $f^{*}=M\mathds{1}_{B(x,R)}$ is the map that maximizes this quantity on $\mathcal{E}_{M,\gamma}$.
The second lemma concerns a variant of Gronwall’s inequality, whose proof is provided in [20, Lemma A.2].
Lemma A.2.
Let $f:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ such that there exists $\kappa$ such that
$\forall\,x,y\in{\mathbb{R}}^{n},\quad\big{|}f(x)-f(y)\big{|}\leqslant\kappa|x-y|.$
Let $g\in L^{1}({\mathbb{R}}_{+},{\mathbb{R}}_{+})$ and $T\geqslant 0$. We assume that $z:{\mathbb{R}}_{+}\to{\mathbb{R}}^{n}$ satisfies
$\forall t\in[0,T],\quad z^{\prime}(t)=f(z(t)),$
that $y:{\mathbb{R}}_{+}\to{\mathbb{R}}^{n}$ satisfies
$\forall t\in[0,T],\quad|y^{\prime}(t)-f(y(t))|\leqslant g(t).$
Then
$\forall t\in[0,T],\quad|y(t)-z(t)|\leqslant\left(\int_{0}^{t}g(s)\mathrm{d}s+|y(0)-z(0)|\right)e^{\kappa t}.$
Proof.
The proof was already done in [20] and is recalled here for the sake of completeness.
We have that for all $t\in[0,T]$,
$|y(t)-z(t)|$ $\leqslant\left|\int_{0}^{t}\big{(}y^{\prime}(s)-z^{\prime}(s)\big{)}\mathrm{d}s\right|+|y(0)-z(0)|$
$\leqslant\int_{0}^{t}g(s)\mathrm{d}s+\left|\int_{0}^{t}\big{(}f(y(s))-f(z(s))\big{)}\mathrm{d}s\right|+|y(0)-z(0)|$
$\leqslant\int_{0}^{t}g(s)\mathrm{d}s+|y(0)-z(0)|+\kappa\int_{0}^{t}|y(s)-z(s)|\mathrm{d}s,$
so using now the classical Gronwall’s inequality, since $t\mapsto\int_{0}^{t}g(s)\mathrm{d}s+|y(0)-z(0)|$ is non negative and differentiable, we have that
$|y(t)-z(t)|\leqslant\left(\int_{0}^{t}g(s)\mathrm{d}s+|y(0)-z(0)|\right)e^{\kappa t}.$
Appendix B On the suitable scaling regime
The purpose of this appendix is to provide a more detailed discussion of suitable scaling regime leading to the leapfrogging dynamics chosen in this paper as well as a detailed comparison to the results of [11]. While the authors obtain a result somewhat in the spirit of Theorem 1.1, the considered asymptotic regime differs significantly for the present one as does the limiting trajectory.
B.1. The scaling regime and respective limiting trajectory
As mentioned in the introduction, a single vortex ring propagates in vertical direction with velocity (1.4) equaling
$v_{L}=|\ln\varepsilon|\frac{\gamma}{4\pi r^{\ast}}e_{z},$
hence depending on the reference radius $r^{\ast}$, the intensity $\gamma$ and in logarithmically diverging manner on its thickness $\varepsilon$. When several vortex rings are considered, their motion is driven by the self-induced motion and their mutual interactions which are known to be of strength that is inversely proportional to their mutual distance. Depending on the chosen asymptotic regime, the former [12, 8] or the latter [10] maybe dominant. The leapfrogging regime corresponds to a regime balancing these two effects being critical in a sense and the rigorous justification of the asymptotic dynamics is significantly more difficult. Indeed, for the 3D Euler equations it remained completely open until the recent aforementioned [18, 11] and our work. As a rule of thumb and outlined in the introduction, the leapfrogging dynamics is expected to occur for vortex rings that travel at leading order at the same self-induced velocity (1.4) of order $|\ln\varepsilon|$, namely share the same reference radius $r^{\ast}$ and intensity $\gamma$. Further, their mutual distance is of order $1/{\sqrt{|\ln\varepsilon|}}$ leading to a mutual interaction of order $\sqrt{|\ln\varepsilon|}$. In a time of order $\mathcal{O}\left(1/|\ln\varepsilon|\right)$ (so of order 1 in the rescaled time variable), this, in turn, generates a variation of the rings’ radii of order $\mathcal{O}\left(1/{\sqrt{|\ln\varepsilon|}}\right)$ which implies a change in their self-induced velocity (1.4) at the order $\sqrt{|\ln\varepsilon|}$, i.e.
$v_{L}=|\ln\varepsilon|\frac{\gamma}{4\pi r^{*}}e_{z}+\mathcal{O}(\sqrt{|\ln\varepsilon|})$
We emphasize that this regime only occurs for rings of the same circulation $\gamma$. If not, the difference of their self-induced velocity (1.4) is of order $|\ln\varepsilon|$, which cannot be balanced by their mutual interaction (of order $\sqrt{|\ln\varepsilon|}$), leading inevitably to the rings moving quickly away from each other. This present scaling regime also corresponds to the one considered in [18, 33]. For sufficiently small $\varepsilon>0$, Theorem 1.1 states that the motion of the vortex rings is governed by (1.8) where $t\mapsto X_{i}^{\varepsilon}(t)$ denotes the trajectory of the $i$-th vortex ring, i.e.
$\begin{cases}\displaystyle\frac{\mathrm{d}}{\mathrm{d}t}X_{i}^{\varepsilon}(t)=\frac{\gamma}{4\pi X_{i,r}^{\varepsilon}(t)}e_{z}+\frac{\gamma}{2\pi|\ln\varepsilon|}\sum_{j\neq i}\frac{\big{(}X_{i}^{\varepsilon}(t)-X_{j}^{\varepsilon}(t)\big{)}^{\perp}}{\big{|}X_{i}^{\varepsilon}(t)-X_{j}^{\varepsilon}(t)\big{|}^{2}},\\ X_{i}^{\varepsilon}(0)=X_{i,0}^{\varepsilon}.\end{cases}$
Moreover, formally assuming that (1.9), namely $X_{i}^{\varepsilon}(t)=X^{*}+\frac{\gamma}{4\pi r^{*}}te_{z}+\frac{1}{\sqrt{|\ln\varepsilon|}}Y_{i}^{\varepsilon}(t)$, holds, we derive the motion law for $Y_{i}^{\varepsilon}(t)$ with $Y_{i}^{\varepsilon}(0)=Y_{i,0}$ obtaining that
$\frac{\mathrm{d}}{\mathrm{d}t}Y_{i}^{\varepsilon}(t)$ $=\frac{\gamma}{2\pi}\sum_{j\neq i}\frac{(Y_{i}^{\varepsilon}-Y_{j}^{\varepsilon})^{\perp}}{|Y_{i}^{\varepsilon}-Y_{j}^{\varepsilon}|^{2}}+\sqrt{|\ln\varepsilon|}\frac{\gamma}{4\pi r^{*}}\left(\frac{1}{1+\frac{Y_{i,r}^{\varepsilon}}{r^{*}\sqrt{|\ln\varepsilon|}}}-1\right)e_{z}$
$=\frac{\gamma}{2\pi}\sum_{j\neq i}\frac{(Y_{i}^{\varepsilon}-Y_{j}^{\varepsilon})^{\perp}}{|Y_{i}^{\varepsilon}-Y_{j}^{\varepsilon}|^{2}}-\frac{\gamma}{4\pi(r^{\ast})^{2}}e_{z}+\mathcal{O}\left(\frac{1}{\sqrt{|\ln\varepsilon|}}\right).$
We recognize that the relative self-induced motion and interaction are of the same order. We refer to Section 5.2 for full details. In particular, it follows from Lemma 5.3 that the limit trajectories $\widetilde{Y}_{i}$ defined in (1.9) satisfy (1.10), i.e.
[…]
We refer to [18, (1.14)] and [33, Section 1.2] and references therein for a qualitative and quantitative analysis of solutions to (B.1) and prove in Appendix C that no collisions occur in finite time.
B.2. Comparison of scaling regimes in the literature
Different from the present the one, see also [18, 35], the authors in [11] consider a scaling regime in which the self-induced motion of the vortex rings is of the same order as the velocity induced by the interaction of vortex rings. We also refer the reader to [42, Section 1.1.1-1.1.2.] for a comparison of both regimes. More precisely, the authors in [11] work with vortex rings of radius $\mathcal{O}(|\ln\varepsilon|)$, vortex circulations of order $\mathcal{O}(1)$ and the vorticity $\Omega$ that reads in terms of the potential vorticity $\omega=\frac{\Omega}{r}$, as follows
$\begin{cases}\partial_{t}\frac{\Omega}{r}+u\cdot\nabla\left(\frac{\Omega}{r}\right)=0,\\ u=\frac{1}{r}\nabla^{\perp}\Psi,\\ \nabla\cdot\left(\frac{1}{r}\nabla\Psi\right)=\Omega\\ \int\Omega_{i}\mathrm{d}x=a_{i},\\ \operatorname{supp}\Omega_{i}\subset B((0,\alpha|\ln\varepsilon|)+x_{i}),\end{cases}$
$\begin{cases}\partial_{t}\omega+u\cdot\nabla\omega=0,\\ u=\frac{1}{r}\nabla^{\perp}\Psi,\\ \nabla\cdot\left(\frac{1}{r}\nabla\Psi\right)=r\omega\\ \int r\omega_{i}\mathrm{d}x=a_{i},\\ \operatorname{supp}\omega_{i}\subset B((0,\alpha|\ln\varepsilon|)+x_{i}).\end{cases}$
For the convenience of the reader, we translate the result of [11] to the scaling regime considered in the present paper. Exploiting the natural scaling invariance of the Euler equations allows one to write this problem in the regime of vortex rings of radius and circulation of order $\mathcal{O}(1)$. To this end, we define
$\widetilde{\omega}(t,x)=|\ln\varepsilon|^{3}\omega(t,|\ln\varepsilon|x),$
and check that the vortex rings’ circulation is order $\mathcal{O}(1)$, specifically
$\int r\widetilde{\omega}_{i}(t,x)\mathrm{d}x=|\ln\varepsilon|^{2}\int(r\omega_{i})(t,|\ln\varepsilon|x)\mathrm{d}x=a_{i}.$
Further, the support satisfies
$\operatorname{supp}\widetilde{\omega}_{i}\subset D\left((0,\alpha)+\frac{x_{i}}{|\ln\varepsilon|}\right),$
so that the scaled rings’ radius of order $\mathcal{O}(1)$. The distances between the rings are now of order $\mathcal{O}(\frac{1}{|\ln\varepsilon|})$, which is more singular than in (1.6) of our analysis. We then compute that
(B.2) $\partial_{t}\widetilde{\omega}(t,x)=|\ln\varepsilon|^{3}\partial_{t}\omega(t,|\ln\varepsilon|x),\quad\nabla\widetilde{\omega}(t,x)=|\ln\varepsilon|^{4}\nabla\omega(t,|\ln\varepsilon|x).$
We then turn to the Biot-Savart law. Our goal is to determine $\widetilde{\Psi}$ such that
$\nabla\cdot\left(\frac{1}{r}\nabla\widetilde{\Psi}\right)=r\widetilde{\omega}.$
Setting
$\widetilde{\Psi}(t,x)=\frac{1}{|\ln\varepsilon|}\Psi(t,|\ln\varepsilon|x),$
one computes that
$\nabla\cdot\left(\frac{1}{r}\nabla\widetilde{\Psi}\right)(t,x)$ $=\frac{1}{|\ln\varepsilon|}\left(-\frac{|\ln\varepsilon|e_{r}}{r^{2}}\cdot\nabla\Psi(t,|\ln\varepsilon|x)+\frac{|\ln\varepsilon|^{2}}{r}\Delta\Psi(t,|\ln\varepsilon|x)\right)$
$=\frac{|\ln\varepsilon|^{3}}{|\ln\varepsilon|}\nabla\cdot\left(\frac{1}{r}\nabla\Psi\right)(t,|\ln\varepsilon|x)$
$=|\ln\varepsilon|^{2}(r\omega)(t,|\ln\varepsilon|x)$
$=|\ln\varepsilon|^{3}r\omega(t,|\ln\varepsilon|x)$
$=r\widetilde{\omega}(t,x),$
so we found the right candidate for $\widetilde{\Psi}$ which leads to
(B.3) $\widetilde{u}(t,x)=\frac{1}{r}\nabla^{\perp}\widetilde{\Psi}(t,x)=\frac{1}{r}\frac{|\ln\varepsilon|}{|\ln\varepsilon|}\nabla^{\perp}\Psi(t,|\ln\varepsilon|x)=|\ln\varepsilon|u(t,|\ln\varepsilon|x).$
Gathering relations (B.2)and (B.3) we get from the equation for the potential vorticity $\omega$ that
$\frac{1}{|\ln\varepsilon|^{3}}\partial_{t}\widetilde{\omega}+\frac{1}{|\ln\varepsilon|^{5}}\widetilde{u}\cdot\nabla\widetilde{\omega}=0$
namely
$\partial_{t}\widetilde{\omega}+\frac{1}{|\ln\varepsilon|^{2}}\widetilde{u}\cdot\nabla\widetilde{\omega}=0.$
In conclusion, the results of [11] can be interpreted as strong confinement result (in both directions) for vortex rings of radius of order $\mathcal{O}(1)$ and circulations of order $\mathcal{O}(1)$, placed at a distance of order $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$, for a time of order $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|^{2}}\right)$. Indeed, one can perform a final change of the time variable
$\hat{\omega}(t,x)=\widetilde{\omega}(t|\ln\varepsilon|^{2},x)\,\qquad\hat{u}(t,x)=\widetilde{u}(t|\ln\varepsilon|^{2},x)\,,$
Finally, the scaled potential vorticity satisfies the transport equation
$\partial_{t}\hat{\omega}+\hat{u}\cdot\nabla\hat{\omega}=0.$
This regime corresponds to the second regime in [42, Section 1.1.2.] where it is expected that the two rings have the same local velocities. The velocities $v_{L}$ (local induction approximation) and $F$ (generated by the other rings) are both of order $|\ln\varepsilon|$. This means that in this regime the vortex rings perform a rotation around each other over a distance of order $\mathcal{O}\left(\frac{1}{|\ln\varepsilon}\right)$ in a time of order $1/|\ln\varepsilon|^{2}$. Moreover, in that regime, the vortex rings can have different circulations $\gamma_{i}$, and in that case the limiting trajectories were proven [11, Theorem 2.1] to satisfy
[…]
Note that (B.4) differs from (B.1) in terms of the vertical translation. The properties of solutions to (B.4) have been investigated in [39], in particular ruling out blow-up provided that all $\gamma_{i}$ are of the same sign [39, Theorem 2.1.]. See also [11, Section 7] for the case of two vortex rings.
For the convenience of the reader we provide a comparison of the choice of parameters between our regime and the one studied in [11].
radius $r^{\ast}$ intensity $\gamma$ velocity $v_{L}$ mutual distance interaction time-scale
our scaling $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(|\ln\varepsilon|)$ $\mathcal{O}\left(\frac{1}{\sqrt{|\ln\varepsilon|}}\right)$ $\mathcal{O}(\sqrt{|\ln\varepsilon|})$ $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|}\right)$
re-scaling of [11] $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(|\ln\varepsilon|)$ $\mathcal{O}\left(\frac{1}{{|\ln\varepsilon|}}\right)$ $\mathcal{O}({|\ln\varepsilon|})$ $\mathcal{O}\left(\frac{1}{|\ln\varepsilon|^{2}}\right)$
Appendix C No collision in finite time
In this section we prove that no collisions in finite time may occur for system (1.10), rewritten here as
[…]
Although this fact is probably well-known, we provide the proof here since we did not find it in the literature.
As observed in [18], the system (C.1) is a Hamiltonian system with Hamiltonian given by
$H_{0}=\frac{1}{2}\sum_{i\neq j}\ln|q_{i}(t)-q_{j}(t)|+\frac{1}{4(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)^{2}.$
Moreover, we observe that the angular momentum
$I(t)=\frac{1}{2}\sum_{i=1}^{N}|q_{i}(t)|^{2}$
satisfies $I(t)\leqslant I(0)e^{\gamma t/4\pi(r^{\ast})^{2}}$.
Indeed, we compute
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}I(t)&=\frac{\gamma}{2\pi}\sum_{i\neq j}q_{i}(t)\cdot\frac{(q_{i}(t)-q_{j}(t))^{\perp}}{|q_{i}(t)-q_{j}(t)|^{2}}-\frac{\gamma}{4\pi(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)q_{i,z}(t)\\ &=-\frac{\gamma}{4\pi(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)q_{i,z}(t)\leqslant\frac{\gamma}{4\pi(r^{\ast})^{2}}I(t),\end{split}$
where we have exchanged $i$ and $j$ in the first sum.
As a consequence, we get by standard arguments
$\begin{split}\sum_{i\neq j||q_{i}-q_{j}|<1}|\ln|q_{i}(t)-q_{j}(t)||&=\sum_{i\neq j||q_{i}-q_{j}|>1}\ln|q_{i}(t)-q_{j}(t)|-2H_{0}+\frac{1}{2(r^{\ast})^{2}}\sum_{i=1}^{N}q_{i,r}(t)^{2}\\ &\leqslant 2\sqrt{2N}\sqrt{I(t)}-2H_{0}+\frac{1}{(r^{\ast})^{2}}I(t).\end{split}$
for all $t\in[0,T^{\ast})$, where $T^{\ast}$ is the maximum time of existence of (C.1), where we have used that for $|x-y|>1$ we have $\ln|x-y|\leqslant|x|+|y|$. The conclusion then follows from the bound on $I(t)$.